### v1

# Latent Assistance Networks: Rediscovering Hyperbolic Tangents in RL

Abstract
Activation functions are one of the key components of a neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and linear-unit functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions. From the perspective of the activations in the last hidden layer, this paper provides insights regarding this sub-optimality and explores how activation functions influence the occurrence of dead neurons and the magnitude of the effective rank. Additionally, a novel neural architecture is proposed that leverages the product of independent activation values. In the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
1 Introduction
Figure 1: Normalized performance training DQN in the Atari domain where the activation of the representation, defined as the final hidden layer, is either ReLU or tanh.
Out of all activation functions, the Rectified Linear Unit (ReLU) (Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016). The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Figure 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019). As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network (Dubey et al., 2022).
In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function remains the most popular activation for performance reasons (Henderson et al., 2018). As a result, continuously differentiable activation functions like the hyperbolic tangent are rarely used (see Fig. 1), even though their symmetrical, bounded shape and smooth gradient landscape has advantages for optimization. Not much attention has been given to investigate and provide a solution to this mysterious lack of performance, despite evidence showing that a hidden layer activated by a hyperbolic tangent is potentially a strong choice, as it leads to a high effective rank and thus network expressivity (Kumar et al., 2021; Gulcehre et al., 2022).
From the perspective of a compressed representation, defined as the activations of the final hidden layer, we therefore investigate the use and improvement of continuously differentiable activations in the reinforcement learning context. This is done by providing an alternative to the conventional way of parameterizing a representation. Specifically, the following contributions are made to the field:
  • We show that, similar to dying ReLU’s, dying hyperbolic tangents are also abundant in the representation when training RL.
  • A Latent Assistance Network (LaN) is introduced, modifying the representation of a network to be the product of two separate, individual activations.
  • We demonstrate that, without tuning any hyperparameters or using any auxiliary losses, the addition of a LaN yields significant performance gains over using a standard continuously differentiable activation in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
2 Preliminaries
Figure 2: A regression of three shallow neural network architectures on a complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$). The Tanh (LaN) network emerges as the strongest function approximator, even while having less trainable parameters (501 vs 601 for Tanh & ReLU). To make a fair comparison, the Tanh and ReLU networks have one single hidden layer of size 200, while the Tanh (LaN) network has one single hidden layer of size 100. For the Tanh (LaN) network however, we use two parallel linear layers preceding the hidden layer in order to be able to use the single hidden layer as the product of two activations (see Section 3). For experiments comparing deeper networks, we refer the reader to Appendix C.1.
We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward mapping and $\gamma$ is the discount factor. A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$. One entry in $B$ contains a tuple of past experience $(s_{t},a_{t},r_{t},s_{t+1})$. The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
Furthermore, we examine the setting where a high-dimensional state ($s_{t}\in\mathbb{R}^{v}$) is compressed into lower-dimensional activations $z_{t}\in\mathcal{Z}=\mathbb{R}^{w}$ where we call $\mathcal{Z}$ the representation space with $w\leq v$. This is done by means of a neural network encoding $e:\mathcal{S}\rightarrow\mathcal{Z}$ where $e$ represents the encoder.
3 Method
Our method aims at using the symmetrical, stable gradient landscape and the naturally bounded properties of a continuously differentiable activation, while avoiding the pitfalls such as the well-known vanishing gradient problem. To accomplish this, we complement the original linear layer with an additional, independent linear layer. Subsequently, we define the final representation as a product of the two linear layers’ activations. In this section, the continuously differentiable activations as well as the full architecture used in our research are explained.
3.1 Continuously Differentiable Activations
Continuously differentiable activations such as the hyperbolic tangent ($tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the $ReLU$ or its Linear-Unit descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity). The hyperbolic tangent and the sigmoid output values in the ranges $[-1,1]$ and $[0,1]$, respectively. These functions are defined as:
$tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}},\qquad\sigma(x)=\frac{1}{1+e^{-x}}$ (1)
Both functions have the advantage of being differentiable everywhere, as well as being bounded. Furthermore, the sigmoid is well suited for output probabilities, while the tanh is convenient when requiring a zero-centered symmetrical output. However, both functions exhibit the vanishing gradient problem for saturating activations (Glorot & Bengio, 2010; Goodfellow et al., 2016).
3.2 Latent assistance Network (LaN)
Figure 3: A visualisation of the Latent assistance Network (LaN) architecture combined with the nature DQN architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment (Mnih et al., 2015). A parallel linear layer providing independent hidden layer activations is integrated, where the element-wise product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
To mitigate saturating activations, an augmentation of the conventional representation architecture is proposed. The networks’ original hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{1}$ and $B_{1}$ representing the layer parameters, and with the function $f()$ that represents a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the number of neurons that are fully saturated (same value for all inputs, similar to collapse), the Latent assistance Network augments the original representation by a parallel representation layer $z^{*}$ that can be seen as a single highway layer with a closed carry gate (Srivastava et al., 2015). The final representation is defined as the element-wise product between the aforementioned activations:
$z(x)=z^{enc}(x)\cdot z^{*}(x)$ (2)
where $z^{*}(x)=f(A_{2}x+B_{2})$ has $A_{2}$ and $B_{2}$ that represent the weight matrix and bias. A visualization of the proposed architecture can be found in Fig. 3.
Vanishing Gradients
Our key hypothesis is that the Latent assistance Network can prevent neurons to saturate easily, hence mitigating vanishing gradients. To explain this phenomenon, we investigate the derivative of a product of two functions. For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $g^{\prime}(x)h(x)+g(x)h^{\prime}(x)$. In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
$z^{\prime}(x)=A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})$
$+A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))$
If a neuron from $f(A_{1}x+B_{1})=0\hskip 2.84526pt\forall\hskip 2.84526ptx$, the gradient of the product becomes 0 while if a positive saturation is experienced i.e. $f(A_{1}x+B_{1})=1\hskip 2.84526pt\forall\hskip 2.84526ptx$, $z^{\prime}(x)$ can remain nonzero. For a product of two hyperbolic tangent functions, the derivative is defined as:
$h^{\prime}(x)=A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})+\tanh(A_{1}x+B_{1})A_{2}\text{sech}^{2}(A_{2}x+B_{2})$
where $\operatorname{sech}^{2}$ is the derivate of a hyperbolic tangent. When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values. This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ can still maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients.
Taking a more formal approximation of neuron collapse through the lens of probability theory, we define the probability of a single neuron saturating as $\mathbf{p}$. Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$. Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron activation function. For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.2.
Hyperbolic Tangent
In the case of the hyperbolic tangent, product saturation only occurs if strictly both neurons are saturated. This results in a probability of $\mathbf{p}\cdot\mathbf{p}=\mathbf{p}^{2}$. Taking a product of hyperbolic tangent activated neurons thus reduces the probability of neuron saturation from $\mathbf{p}$ to $\mathbf{p}^{2}$.
Sigmoid
For the sigmoid function, product saturation occurs in two scenarios: Either one of the neurons is saturated towards zero or both neurons are saturated towards 1. The probability that a single neuron does not saturate towards zero is $(1-0.5\mathbf{p})$, and subsequently the probability that neither neuron saturates towards zero is $(1-0.5\mathbf{p})^{2}$. The probability that at least one of the two neurons saturates to zero is therefore $1-(1-0.5\mathbf{p})^{2}=\mathbf{p}-0.25\mathbf{p}^{2}$. Adding the probability that both neurons saturate towards 1, which is $(0.5\mathbf{p})^{2}=0.25\mathbf{p}^{2}$, the final probability of the neuron product saturation is $\mathbf{p}-0.25\mathbf{p}^{2}+0.25\mathbf{p}^{2}=\mathbf{p}$. Taking a product of sigmoid activated neurons therefore does not reduce the probability of neuron collapse.
ReLU
In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$. As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$. The probability that at least one neuron saturates is thus equal to $1-(1-\mathbf{p})^{2}=2\mathbf{p}-\mathbf{p}^{2}$. As the ReLU saturation results in strict zeroes, this results in the product also being zero. Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$.
Table 1: Summary of dying neuron probabilities with and without a Latent assistance Network (LaN).
Activation Function Dying Neuron Probability Probability with LaN Difference
Hyperbolic Tangent $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
Sigmoid $\mathbf{p}$ $\mathbf{p}$ 0
ReLU $\mathbf{p}$ $2\mathbf{p}-\mathbf{p}^{2}$ $+(\mathbf{p}-\mathbf{p}^{2})$
It might seem logical to address neuron saturation by increasing the representation layer size, e.g. doubling it. However, in the experiments section we show that larger layers are not desirable and fail to achieve the unique benefits of the introduced architecture.
Furthermore, the product function $h(x)g(x)$ can exhibit more complex behavior due to the interactions between the input parameters influenced by both functions. These properties are illustrated with a basic regression task in Fig. 2.
(a) Tanh - $10^{6}$ iterations
(b) Tanh - $10^{7}$ iterations
(c) Tanh (LaN) - $10^{6}$ iterations
(d) Tanh (LaN) - $10^{7}$ iterations
Figure 4: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the ’Breakout’ environment. Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$. The representation using a LaN tends to more quickly utilize the full range of the hyperbolic tangent and strongly mitigates dying neurons as compared to using a single hyperbolic tangent. Furthermore, using a LaN leads to more multimodal density distribution as compared to using a regular hyperbolic tangent. More information on the neuron KDE’s can be found in Appendix B.
Dying Hyperbolic Tangents
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that representations activated by hyperbolic tangents tend to also show strong dying neuron behavior. Given that the hyperbolic tangent is an asymptotic function near its saturation point, the classification of its saturation remains qualitative. In order to combine visualization and neuron classification, we approximate the amount of dying hyperbolic tangents by applying a Kernel Density Estimation (KDE) to the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the compressed representation $z_{t}$. More details of the KDE calculation and neuron classification can be found in Appendix B. For the visualization of the representation layer, a fixed subset of the neurons $\alpha_{i}$ is taken. For both an encoder with and without using a LaN, we provide a visualization of sixteen individual neuron KDE’s in Fig. 4.
4 Experiments
We conduct our main experiments and discussion on the Deep Q-Learning algorithm (DQN) by Mnih et al. (2015) in Atari, and complement them by additionally evaluating the LaN on the Proximal Policy Optimization algorithm (PPO) (Schulman et al., 2017). In all experiments, we build on the strong baseline implementations from cleanrl  (Huang et al., 2022). For both DQN and PPO, we define the representation $z_{t}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig. 3). Using a LaN, we show that it is possible to decrease dying hyperbolic tangents while respectively increasing the effective rank of the representation $z_{t}$, as well as downstream performance.
(a) Normalized Score
(b) Ablations
Figure 5: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames). In (a), a product of hyperbolic tangents is the only continuously differentiable activation that performs equal or better than the ReLU activation, without any hyperparameter changes. Furthermore, using a product of hyperbolic tangent representations seems to greatly outperform the single hyperbolic tangent. In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown. Notably, the decrease in performance when using a single hyperbolic tangent with a twice as large representation ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable.
We perform all experiments on 8 common, non-exploration driven Atari games. We normalize using the ReLU baseline normalized scores and show in Appendix D.1 that human-normalized scores provide an even better score for Tanh (LaN), due to the strong LaN performance in one of the games where humans perform much worse.
Reinforcement Learning
In DQN, the action $a_{t}$ is chosen following an $\epsilon$-greedy policy. With probability $\epsilon$, a random action is selected, and with $(1-\epsilon)$, the action maximizing the Q-value is chosen. The target $Y_{t}$ is defined as:
$Y_{t}=r_{t}+\gamma Q^{\prime}(z_{t+1},\underset{a\in\mathcal{A}}{\operatorname*{arg\,max}}Q(z_{t+1},a)),$ (3)
where $Q^{\prime}(z,a)$ denotes the target Q-network, an auxiliary network that stabilizes the learning by providing a stable target for $Q(z,a)$. The parameters of $Q^{\prime}$ are updated less frequently to enhance learning stability. The loss function for training the network is:
$\mathcal{L}_{Q}=\big{|}Y_{t}-Q(z_{t},a)\big{|}^{2}.$ (4)
Proximal Policy Optimization (PPO) operates on a different principle, utilizing policy gradient methods for policy improvement. PPO seeks to update the policy by maximizing an objective function while preventing large deviations from the previous policy through a clipping mechanism in the objective’s estimator. The clipped policy gradient loss $L^{CLIP}$ is defined as:
$L^{CLIP}(\theta)=\mathbb{E}\left[\min(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right],$ (5)
where $r_{t}(\theta)$ represents the ratio of the probabilities under the new policy versus the old policy, and $\hat{A}_{t}$ is the advantage estimate at timestep $t$. This clipped surrogate objective ensures gradual and stable policy updates.
4.1 Atari - DQN
(a) Dead Neurons
(b) Effective Rank
Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames). Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior when training RL. Furthermore, a product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
Figure 7: Additional ablations run on DQN in the Atari domain. Tanh (+LaN) represents an addition rather than an elementwise product with a LaN, and Tanh (2xLaN) uses two Latent assistance Networks.
The influence of a LaN on downstream performance can be seen in Fig. 5. A significant improvement over the standard hyperbolic tangent baseline is seen, as well as an important improvement over the strong ReLU baseline. Examining further ablations in Fig. 7 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance. Furthermore, taking a product of 3 hyperbolic tangents (2$\times$×LaN) also appears to enhance performance, though the improvement is not significantly different from using a single LaN. This could be the result of increased contracting behavior in the early stage of training due to subsequent multiplication of activations whose absolute values are $<1$. Further experiments combining the ReLU activation with a LaN can be found in Appendix C.3.
Mitigating dying Hyperbolic Tangents
As mentioned in Section 3, the number of dead neurons is defined as the amount of neurons that display the same output for any given observation $s_{t}$. The amount of dead neurons over time when training on Atari can be seen in  6(a). For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ are dead, while for the sigmoid and hyperbolic tangent activation this number is around 40%. When using a hyperbolic tangent with a LaN, a reduction in dead neurons as compared to using a single hyperbolic tangent or sigmoid can be observed. We credit this to the inherent ability of a product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3. This reduction of long-term dead hyperbolic tangents when using the LaN is remarkable as it is also coinciding with a stronger divergence of values in the initial stages of training (see Fig. 4).
Increasing Effective Rank
We calculate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig. 6(b). As observed by Gulcehre et al. (2022), a representation activated by a hyperbolic tangent or, residing in the same category, a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU. Furthermore, as can additionally be seen in supervised learning (see Fig. 2), a LaN notably improves the effective rank of the hyperbolic tangent, which can be seen as a measure of network ‘expressivity’. We summarize our results qualitatively in Table 2.
Table 2: High-level summary of the strengths of a LaN.
Activation function Performance RL Few dead neurons Effective rank
Tanh (LaN) ++ ++ ++
Tanh - - +
ReLu +
Sigmoid - - -
4.2 Atari - PPO
Additional experiments are run using the PPO algorithm (Schulman et al., 2017). For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2). The results can be found in Fig. 8. Similarly to DQN, the Tanh (LaN) exhibits the highest effective rank and the best performance. The Tanh (LaN) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations are relatively unreliable.
(a) Normalized Score
(b) Effective Rank
Figure 8: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
5 Related Work
Network Capacity in RL.
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank and do not suffer strongly from rank decay as training continues. Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Arnob et al., 2021). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
Network Architecture
The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject. Work by Srivastava et al. (2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997). Using these ideas, the Resnet was invented (He et al., 2016). More specific to RL, Henderson et al. (2018) showed differences in RL performances over different network architectures and nonlinear activations. Work by Abbas et al. (2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch. Finally, recent work by Grooten et al. (2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
6 Conclusions and Discussion
We have proposed a novel architecture called the Latent assistance Network (LaN) that augments a networks’ final hidden layer by taking the element-wise product with a parallel, independent hidden layer. We discussed and empirically showed that this decreases the occurrence of dead neurons in the representation $z_{t}$, increases network expressiveness, and notably improves performance compared to a single continuously differentiable activation such as the hyperbolic tangent.
Our results also show an important increase in performance over the dominant ReLU activation. This paper therefore remains, to the best of our knowledge, the first to show the possibility of elevating a continuously differentiable activation to and beyond the level of the ReLU. Additionally, it should be recognized that the employed DQN baseline was initially optimized specifically for the ReLU activation function.
As investigated by Kumar et al. (2021), a compressed representation activated by a hyperbolic tangent has a much higher effective rank than the ReLU and could therefore yield stronger network expressiveness. We therefore remain hopeful that the hyperbolic tangent can potentially replace the ReLU in a variety of architectures. Making strong claims in this field however remains challenging, as the success of pruning strategies show that network expressiveness might already be in abundance in some settings (Frankle & Carbin, 2019; Obando-Ceron et al., 2024).
Future work could focus on further identifying the intricacies of this architecture in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in supervised learning. Also, we believe that an implementation of this architecture in a continual learning setting with similar evaluations as the work in Abbas et al. (2023) could be promising.
7 Limitations
Our work has focused on the encoders’ final hidden layer in reinforcement learning. As literature suggests that convolutional layers have an even stronger preference for ReLU activations (Jarrett et al., 2009), we have not conducted experiments testing a LaN on the convolutional activations of our encoders.
Due to the heavy computational requirements of the Atari domain, limitations of our work include the use of a moderate amount of 5 seeds per environment, as well as the lack of a hyperparameter search for using the ‘Tanh (LaN)’ setting in our experiment section. A hyperparameter search could further strengthen our results, as the strong baseline is specifically tuned for the ReLU activation. Furthermore, as we perform experiments on DQN and PPO, integration of our Latent assistance Network into more sophisticated algorithms such as Rainbow (Hessel et al., 2018) would be interesting.
Appendix A Implementation Details
A.1 Hyperparameters
To evaluate, 8 different Atari environments are tested, using 5 different random seeds. For the mean scores, we take the mean over the eight environments. Our normalized score is calculated according to our baseline, the original implementation using a ReLU activation.
All the hyperparameters used in our experiments for DQN and PPO, respectively, are as reported in cleanrl (Huang et al., 2022). The hyperparameters can be found in Table 1 and Table 2.
Table 3: DQN Hyperparameters
Hyperparameter Value Description
Learning Rate $1\times 10^{-4}$ Learning rate for the optimizer
Discount Factor $(\gamma)$ 0.99 Discount for future rewards
Replay Memory Size 1,000,000 Size of the experience replay buffer
Mini-batch Size 32 Number of samples per batch update
Target Network Update Frequency 1000 Update frequency for the target network
Initial Exploration 1.0 Initial exploration rate in $\epsilon$-greedy
Final Exploration 0.1 Final exploration rate in $\epsilon$-greedy
Final Exploration Frame 1,000,000 Frame number to reach final exploration
Exploration Decay Frame 1,000,000 Frames over which exploration rate decays
Action Repeat (Frame Skip) 4 Number of frames skipped per action
Reward Clipping [-1, 1] Range to which rewards are clipped
Input Dimension 84 x 84 Dimensions of the input image
Latent Dimension 512 Dimension of the latent representation
Input Frames 4 Number of frames used as input
Training Start Frame 80,000 Frame number to start training
Loss Function Mean Squared Error Loss function used for updates
Optimizer Adam Optimization algorithm used
Optimizer $\epsilon$ $10^{-5}$ Adam Epsilon
Table 4: PPO Hyperparameters
Hyperparameter Value Description
Learning Rate $2.5\times 10^{-4}$ Learning rate for the optimizer
Discount Factor $(\gamma)$ 0.99 Discount factor for future rewards
Number of Steps 128 Number of steps per environment before update
Anneal LR True Whether to anneal the learning rate
GAE Lambda 0.95 Lambda parameter for GAE
Number of Minibatches 4 Number of minibatches to split the data
Update Epochs 4 Number of epochs per update
Normalize Advantage True Whether to normalize advantage estimates
Clipping Coefficient 0.1 Clipping parameter for PPO
Clip Value Loss True Whether to clip value loss
Entropy Coefficient 0.01 Coefficient for entropy bonus
Value Function Coefficient 0.5 Coefficient for value function loss
Maximum Gradient Norm 0.5 Maximum norm for gradient clipping
Target KL None Target KL divergence between updates
Latent Dimension 512 Dimension of the latent representation
Optimizer Adam Optimization algorithm used
Optimizer $\epsilon$ $10^{-5}$ Adam Epsilon
A.2 PPO Architecture
Figure 9: A visualisation of the Latent assistance Network (LaN) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment. A parallel linear layer providing an independent representation is integrated, where the element-wise product of the two parallel representations represents the final representation $z_{t}$.
Appendix B Kernel Density Estimations
As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without a LaN are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons. We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig. 10(a).
(a) Tanh (LaN) - $5\cdot 10^{6}$ iterations
(b) Tanh (LaN) - $5\cdot 10^{6}$ iterations
(c) Tanh - $5\cdot 10^{6}$ iterations
Figure 10: Kernel Density Estimations (KDE) over a subset of 16 neurons in the representations $z_{t}^{enc}$ and $z^{*}_{t}$ in (a), the resulting element-wise product $z_{t}$ in (b) and the representation $z_{t}$ when training without a LaN (c). These representations are obtained after training DQN in the ’Breakout’ environment. A closer look at neurons 3, 8 and 9 show that when one of the representations saturates, the other is able to compensate, leading to a non-dead neuron in their product $z_{t}$.
(a) Tanh - $10^{7}$ iterations
(b) Tanh (LaN) - $10^{7}$ iterations
(c) Sigmoid - $10^{7}$ iterations
(d) Sigmoid (LaN) - $10^{7}$ iterations
Figure 11: Kernel Density Estimations of the final representation $z_{t}$ after training DQN for $10^{7}$ iterations in the Breakout environment. A hyperbolic tangent using LaN allows the representation to avoid strong saturation, keeping sufficient kernel density in the central sections of the hyperbolic tangent. As a sigmoid can saturate into zero, using a LaN remains less effective into preventing saturation.
B.1 KDE calculation
Firstly, to stabilize the KDE computation and avoid singularity issues, a small noise $\epsilon$, following a normal distribution, is added to each neuron’s activations:
$\alpha^{\prime}_{i}=\alpha_{i}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$
where $\sigma^{2}=1\times 10^{-5}$. The bandwidth for KDE, crucial for the accuracy of the density estimate, is calculated using Scott’s rule, adjusted by the standard deviation of the jittered activations:
$bw=n^{-\frac{1}{5}}\cdot\text{std}(\alpha^{\prime}_{i})$
where $n$ is the number of samples in $\alpha_{i}$. The density of activations is then estimated using a Gaussian kernel:
$f(x)=\frac{1}{n\cdot bw}\sum_{j=1}^{n}K\left(\frac{x-\alpha^{\prime}_{ij}}{bw}\right)$
Here, $K$ denotes the Gaussian kernel function. In order to finally determine if a neuron is dead, the maximum value of the estimated density function $f(x)$ is compared against a predefined threshold:
$\max(f(x))\geq\omega$
where $\omega$ represents the predetermined threshold. In practice, after analyzing the individual neuron KDE’s, using an $\omega$ of 20 provides a strong approximation of actual dead neurons.
Appendix C Additional Experiments
C.1 Shallow and Deep Function Approximation
To further showcase the effect of activations on complex function approximation, we compare the single hidden layer Tanh (LaN) network from Fig. 2 with a deep ReLU and Tanh network containing three hidden layers each. The comparison with shallow networks can be found in Fig. 12(a) and a comparison with deep networks can be found in Fig. 12(b).
(a) Comparison of shallow networks for a nonlinear regression task. The Tanh and ReLU networks have a single hidden layer of 200 neurons, while the Tanh (LaN) has a single hidden layer of 100 neurons but two preceding linear layers. The Tanh and ReLU networks have 601 parameters, while the Tanh (LaN) network has 501 parameters. As found by Gulcehre et al. (2022), a shallow network activated by ReLU has a lower effective rank and consequently reduced network expressivity as compared to a Tanh activated network. Using a LaN, we achieve better function approximation while using less parameters.
(b) Comparison of two deep networks and one shallow network for the same nonlinear regression task. The Tanh and ReLU networks have 3 hidden layers of 200 neurons each, while the Tanh (LaN) network remains shallow. In line with common observations in deep learning, the ReLU activation thrives in deeper networks, in contrast to the Tanh activation. Interestingly, the shallow Tanh (LaN) network still achieves better function approximation with only 0.6% of the deeper networks’ parameters (81001 vs 501). No hyperparameter tuning or architecture search has been applied. Additional tests using deep Tanh (LaN) networks gave similar function approximation as compared to the shallow Tanh (LaN) network.
C.2 Validation of dying neuron probability derivations
As discussed in Section 3, the effect of adding a Latent assistance Network (LaN) to the original representation strongly depends on the activation function. These derivations are empirically validated by the results in Fig. 13. In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1. This could explain the very slight increase in dead neurons when using a LaN with a sigmoid.
(a) Sigmoid with and without LaN
(b) Tanh with and without LaN
(c) ReLU with and without LaN
Figure 13: By evaluating the effect of a LaN on dying neurons through the lens of probability theory, we predicted that only the hyperbolic tangent benefits in this metric. Specifically, only a hyperbolic tangent was speculated to have a decrease in dying neurons. Using a LaN on a sigmoid activation would have no notable difference, and for the ReLU activation we expected an increase in dead neurons. This empirically validates our theory in Section 3.
C.3 Combining ReLU with a LaN
Additional Atari experiments are provided comparing a ReLU activation with and without a LaN. The results can be seen in fig. 14.
(a) Normalized Score
(b) Dead Neurons
(c) Effective Rank
Figure 14: As a Rectified Linear Unit creates sparse representations, it does not benefit from using a LaN, since the final representation will consist of the element-wise product between two sparse representations. Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
Appendix D Atari
D.1 Evaluation Details
We normalize performance with respect to the ReLU baseline in cleanrl  Huang et al. (2022). The minimum and maximum score of the ReLU baseline are taken for each environment, and the normalized score for each environment is calculated as follows:
Normalized Score=ScoreMin ScoreMax ScoreMin $\text{Normalized Score}=\frac{\text{Score}-\text{Min Score}}{\text{Max Score}-\text{Min Score}}$ (6)
where Score refers to the raw performance score of the model being evaluated, Min Score is a single value representing the lowest score obtained by the ReLU baseline (usually equivalent to random policy or even slightly worse), and Max Score is a single value representing the highest score achieved by the ReLU baseline in the same environment. To average, we sum the normalized scores for every run and take the mean.
The more official Human-Normalized Score, as referenced in Mnih et al. (2015), is calculated similarly but using human and random performance benchmarks:
[…]
where Human Score and Random Score refer to the scores recorded by human players and random agents, respectively. Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig. 15. Due to taking a subset of the full atari domain, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation. For a more realistic comparison of the methods, we therefore decided to use baseline-normalized scores in the main paper.
Figure 15: Human-Normalized performance (in multiples) with the standard deviation over the means in the Atari domain for 10M iterations (40M Frames)
D.2 Individual Environments
Figure 16: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. To avoid cherry-picking, no other Atari games were tested.
Figure 17: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. To avoid cherry-picking, no other Atari games were tested.

### v2

@@
-Latent Assistance Networks: Rediscovering Hyperbolic Tangents in RL
+Hadamard Representations: Augmenting Hyperbolic Tangents in RL
 Abstract
-Activation functions are one of the key components of a neural network.
+Activation functions are one of the key components of a deep neural network.
 The most commonly used activation functions can be classed into the category of continuously differentiable (e.g.
@@
 In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions.
-From the perspective of the activations in the last hidden layer, this paper provides insights regarding this sub-optimality and explores how activation functions influence the occurrence of dead neurons and the magnitude of the effective rank.
-Additionally, a novel neural architecture is proposed that leverages the product of independent activation values.
-In the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
+We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s.
+To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation.
+Using deep Q-networks and proximal policy optimization in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
 1 Introduction
-Figure 1: Normalized performance training DQN in the Atari domain where the activation of the representation, defined as the final hidden layer, is either ReLU or tanh.
-Out of all activation functions, the Rectified Linear Unit (ReLU) (Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016).
+Figure 1: Normalized performance training DQN in the Atari domain where the nonlinear activation function of the representation, defined as the final hidden layer, is either ReLU or tanh.
+A strong performance discrepancy due to a relatively small architectural change can be observed.
+Out of all activation functions, the Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016).
 The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
-A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Figure 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019).
-As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network (Dubey et al., 2022).
-In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function remains the most popular activation for performance reasons (Henderson et al., 2018).
-As a result, continuously differentiable activation functions like the hyperbolic tangent are rarely used (see Fig.
-1), even though their symmetrical, bounded shape and smooth gradient landscape has advantages for optimization.
-Not much attention has been given to investigate and provide a solution to this mysterious lack of performance, despite evidence showing that a hidden layer activated by a hyperbolic tangent is potentially a strong choice, as it leads to a high effective rank and thus network expressivity (Kumar et al., 2021; Gulcehre et al., 2022).
-From the perspective of a compressed representation, defined as the activations of the final hidden layer, we therefore investigate the use and improvement of continuously differentiable activations in the reinforcement learning context.
-This is done by providing an alternative to the conventional way of parameterizing a representation.
-Specifically, the following contributions are made to the field:
-  • We show that, similar to dying ReLU’s, dying hyperbolic tangents are also abundant in the representation when training RL.
-  • A Latent Assistance Network (LaN) is introduced, modifying the representation of a network to be the product of two separate, individual activations.
-  • We demonstrate that, without tuning any hyperparameters or using any auxiliary losses, the addition of a LaN yields significant performance gains over using a standard continuously differentiable activation in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
+A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Fig.
+2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019).
+As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
+In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018).
+Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig.
+1).
+However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks.
+Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022).
+Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
+We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity.
+Based on these insights, we mitigate said effects by augmenting the original architecture.
+Specifically, we focus our research on the activations of the compressed representation of an encoder, defined as the final hidden layer, and provide an alternative to the conventional parameterization of this representation.
+Our contributions can be summarized as follows:
+  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem.
+  • A Hadamard representation (HR) is proposed, defining a latent representation as the Hadamard product of two separate, individually parameterized activation vectors.
+  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, a Hadamard representation yields significant performance gains over a standard hyperbolic tangent in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
 2 Preliminaries
-Figure 2: A regression of three shallow neural network architectures on a complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$).
-The Tanh (LaN) network emerges as the strongest function approximator, even while having less trainable parameters (501 vs 601 for Tanh & ReLU).
-To make a fair comparison, the Tanh and ReLU networks have one single hidden layer of size 200, while the Tanh (LaN) network has one single hidden layer of size 100.
-For the Tanh (LaN) network however, we use two parallel linear layers preceding the hidden layer in order to be able to use the single hidden layer as the product of two activations (see Section 3).
+Figure 2: A regression of three shallow neural network architectures on a random complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$).
+The Tanh (HR) network emerges as the strongest function approximator, even while having less trainable parameters (501 vs 601 for Tanh & ReLU).
+To make a fair comparison, the Tanh and ReLU networks have one single hidden layer of size 200, while the Tanh (HR) network has a hidden layer of size 100.
+For the Tanh (HR) network however, we use two parallel linear layers preceding the hidden layer in order to be able to use the single hidden layer as the Hadamard product of two activations (see Section 3).
 For experiments comparing deeper networks, we refer the reader to Appendix C.1.
 We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$.
-Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward mapping and $\gamma$ is the discount factor.
+$\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor.
 A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$.
@@
 This is done by means of a neural network encoding $e:\mathcal{S}\rightarrow\mathcal{Z}$ where $e$ represents the encoder.
-3 Method
-Our method aims at using the symmetrical, stable gradient landscape and the naturally bounded properties of a continuously differentiable activation, while avoiding the pitfalls such as the well-known vanishing gradient problem.
-To accomplish this, we complement the original linear layer with an additional, independent linear layer.
-Subsequently, we define the final representation as a product of the two linear layers’ activations.
-In this section, the continuously differentiable activations as well as the full architecture used in our research are explained.
-3.1 Continuously Differentiable Activations
+3 Augmenting Hyperbolic Tangents
 Continuously differentiable activations such as the hyperbolic tangent ($tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the $ReLU$ or its Linear-Unit descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity).
@@
 However, both functions exhibit the vanishing gradient problem for saturating activations (Glorot & Bengio, 2010; Goodfellow et al., 2016).
-3.2 Latent assistance Network (LaN)
-Figure 3: A visualisation of the Latent assistance Network (LaN) architecture combined with the nature DQN architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment (Mnih et al., 2015).
-A parallel linear layer providing independent hidden layer activations is integrated, where the element-wise product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
-To mitigate saturating activations, an augmentation of the conventional representation architecture is proposed.
-The networks’ original hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{1}$ and $B_{1}$ representing the layer parameters, and with the function $f()$ that represents a nonlinear activation function while $x$ is the set of activations from the previous layer.
-In order to reduce the number of neurons that are fully saturated (same value for all inputs, similar to collapse), the Latent assistance Network augments the original representation by a parallel representation layer $z^{*}$ that can be seen as a single highway layer with a closed carry gate (Srivastava et al., 2015).
-The final representation is defined as the element-wise product between the aforementioned activations:
-$z(x)=z^{enc}(x)\cdot z^{*}(x)$ (2)
-where $z^{*}(x)=f(A_{2}x+B_{2})$ has $A_{2}$ and $B_{2}$ that represent the weight matrix and bias. A visualization of the proposed architecture can be found in Fig. 3.
-Vanishing Gradients
-Our key hypothesis is that the Latent assistance Network can prevent neurons to saturate easily, hence mitigating vanishing gradients.
-To explain this phenomenon, we investigate the derivative of a product of two functions.
-For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $g^{\prime}(x)h(x)+g(x)h^{\prime}(x)$.
+Dying Hyperbolic Tangents
+Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior.
+When using a hyperbolic tangent in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, can be classified as dying if:
+$|\alpha_{i}|\approx 1\;\forall\;s_{t}\in B$ (2)
+Where $s_{t}$ is an observation in buffer $B$.
+Note that an absolute value is used, as a hyperbolic tangent can either saturate towards -1 or 1.
+Given that the hyperbolic tangent is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) .
+To approximate the condition given in Eq.
+2, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer.
+In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken.
+A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig.
+3.
+(a) Tanh - $10^{6}$ iterations
+(b) Tanh - $10^{7}$ iterations
+Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN Mnih et al.
+(2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$.
+Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
+Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
+More details on the KDE calculation and dying neuron classification can be found in Appendix B.
+Hadamard Representations (HR)
+As Fig.
+3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed.
+In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer.
+In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$.
+This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015).
+The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$.
+A visualization of the proposed architecture can be found in Fig.
+4.
+Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015).
+Vertical and horizontal bars represent convolutional and linear layers, respectively.
+A parallel linear layer providing independently parameterized hidden layer activations is integrated, where the Hadamard product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
+Preventing Dying Neurons
+Our key hypothesis is that the Hadamard representation can prevent saturation, hence alleviating vanishing gradients and dying neurons (See Eq.
+2).
+To support our hypothesis, we investigate the derivative of a product of two functions.
+For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}g(x)h^{\prime}(x)}}$.
 In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
-$z^{\prime}(x)=A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})$
-$+A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))$
+$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}$
+$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
 If a neuron from $f(A_{1}x+B_{1})=0\hskip 2.84526pt\forall\hskip 2.84526ptx$, the gradient of the product becomes 0 while if a positive saturation is experienced i.e.
@@
 For a product of two hyperbolic tangent functions, the derivative is defined as:
-$h^{\prime}(x)=A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})+\tanh(A_{1}x+B_{1})A_{2}\text{sech}^{2}(A_{2}x+B_{2})$
-where $\operatorname{sech}^{2}$ is the derivate of a hyperbolic tangent.
+$h^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
+where $\operatorname{sech}^{2}$ is the derivative of a hyperbolic tangent.
 When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values.
-This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ can still maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients.
-Taking a more formal approximation of neuron collapse through the lens of probability theory, we define the probability of a single neuron saturating as $\mathbf{p}$.
+This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ is still able to maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients.
+We visualize the kernel densities during training with a Hadamard representation in Fig.
+5.
+The individual representations before taking the Hadamard product can be found in Appendix B.
+Taking a more formal approximation of neuron collapse, we start by defining the probability of a single neuron saturating as $\mathbf{p}$.
 Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$.
-Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron activation function.
-For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.2.
-Hyperbolic Tangent
+Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron’s activation function.
+(a) Tanh (HR) - $10^{6}$ iterations
+(b) Tanh (HR) - $10^{7}$ iterations
+Figure 5: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the Breakout environment using a Hadamard representation (HR) with hyperbolic tangents.
+The Hadamard representation tends to quickly utilize the full range of the hyperbolic tangent while also mitigating dying neurons.
+Hyperbolic Tangent:
 In the case of the hyperbolic tangent, product saturation only occurs if strictly both neurons are saturated.
@@
 Taking a product of hyperbolic tangent activated neurons thus reduces the probability of neuron saturation from $\mathbf{p}$ to $\mathbf{p}^{2}$.
-Sigmoid
+Sigmoid:
 For the sigmoid function, product saturation occurs in two scenarios: Either one of the neurons is saturated towards zero or both neurons are saturated towards 1.
@@
 Taking a product of sigmoid activated neurons therefore does not reduce the probability of neuron collapse.
-ReLU
+Rectified Linear-Unit:
 In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$.
@@
 Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$.
-Table 1: Summary of dying neuron probabilities with and without a Latent assistance Network (LaN).
-Activation Function Dying Neuron Probability Probability with LaN Difference
+For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.3.
+Table 1: Predicted dying neuron probabilities with and without a Hadamard representation.
+Activation Function Dying Neuron Probability Probability with HR Difference
 Hyperbolic Tangent $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
@@
 ReLU $\mathbf{p}$ $2\mathbf{p}-\mathbf{p}^{2}$ $+(\mathbf{p}-\mathbf{p}^{2})$
-It might seem logical to address neuron saturation by increasing the representation layer size, e.g.
-doubling it.
-However, in the experiments section we show that larger layers are not desirable and fail to achieve the unique benefits of the introduced architecture.
-Furthermore, the product function $h(x)g(x)$ can exhibit more complex behavior due to the interactions between the input parameters influenced by both functions.
-These properties are illustrated with a basic regression task in Fig.
-2.
-(a) Tanh - $10^{6}$ iterations
-(b) Tanh - $10^{7}$ iterations
-(c) Tanh (LaN) - $10^{6}$ iterations
-(d) Tanh (LaN) - $10^{7}$ iterations
-Figure 4: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the ’Breakout’ environment.
-Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
-The representation using a LaN tends to more quickly utilize the full range of the hyperbolic tangent and strongly mitigates dying neurons as compared to using a single hyperbolic tangent.
-Furthermore, using a LaN leads to more multimodal density distribution as compared to using a regular hyperbolic tangent.
-More information on the neuron KDE’s can be found in Appendix B.
-Dying Hyperbolic Tangents
-Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that representations activated by hyperbolic tangents tend to also show strong dying neuron behavior.
-Given that the hyperbolic tangent is an asymptotic function near its saturation point, the classification of its saturation remains qualitative.
-In order to combine visualization and neuron classification, we approximate the amount of dying hyperbolic tangents by applying a Kernel Density Estimation (KDE) to the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the compressed representation $z_{t}$.
-More details of the KDE calculation and neuron classification can be found in Appendix B.
-For the visualization of the representation layer, a fixed subset of the neurons $\alpha_{i}$ is taken.
-For both an encoder with and without using a LaN, we provide a visualization of sixteen individual neuron KDE’s in Fig.
-4.
 4 Experiments
-We conduct our main experiments and discussion on the Deep Q-Learning algorithm (DQN) by Mnih et al.
-(2015) in Atari, and complement them by additionally evaluating the LaN on the Proximal Policy Optimization algorithm (PPO) (Schulman et al., 2017).
-In all experiments, we build on the strong baseline implementations from cleanrl  (Huang et al., 2022).
-For both DQN and PPO, we define the representation $z_{t}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig.
-3).
-Using a LaN, we show that it is possible to decrease dying hyperbolic tangents while respectively increasing the effective rank of the representation $z_{t}$, as well as downstream performance.
+The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017).
+In all experiments, we build on the implementations from cleanrl  (Huang et al., 2022).
+For more details on the PPO and DQN implementation, we refer the reader to Appendix A.1.
+For both DQN and PPO, we define the representation $z_{t}\in\mathbb{R}^{512}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig.
+4).
+We perform all experiments on 8 common, non-exploration driven Atari games.
+We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance.
+For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix.
+B.1 and Appendix.
+B.2, respectively.
+The performance is aggregated over 8 environments and normalized using the ReLU baseline scores.
+(a) Dead Neurons
+(b) Effective Rank
+Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames).
+Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior during training.
+A Hadamard product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
+Mitigating dying Hyperbolic Tangents
+As mentioned in Section 3 and defined in Eq.
+2, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$.
+The amount of dead neurons over time when training on Atari can be seen in Fig.
+6(a).
+For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ die during training, while for the sigmoid and hyperbolic tangent activation this number is around 40%.
+When using a hyperbolic tangent Hadamard representation, a reduction in dead neurons as compared to using a single hyperbolic tangent can be observed.
+We credit this to the inherent ability of a Hadamard product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3.
+Quantitative results of dead neurons can be found in Table 2, which is correlated with earlier predictions from Table 1.
+Table 2: Dying neuron fractions with and without a Hadamard representation (HR).
+Activation Function Dead Neuron Fraction Dead Neuron Fraction (HR) Difference
+Hyperbolic Tangent 0.39 0.30 -23%
+Sigmoid 0.44 0.45 +2%
+ReLU 0.62 0.73 +18%
+For more dying neuron graphs related to the ReLU and the Sigmoid activations, we refer the reader to Appendix C.3.
+Increasing Effective Rank
+We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig.
+6(b).
+As observed by Gulcehre et al.
+(2022), a representation activated by a hyperbolic tangent or a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU.
+Furthermore, similar to the results in our supervised learning experiments (see Fig.
+2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’.
+Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected.
+For more information on the effect of a Hadamard representation on a ReLU, we refer the reader to Appendix C.4.
+Analyzing Performance in Atari
+The influence of a Hadamard representation on downstream performance can be seen in Fig.
+7.
+Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement of nearly $60\%$ over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline.
+Note that no hyperparameter or architectural tuning has been employed in our experiments.
 (a) Normalized Score
@@
 Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames).
-In (a), a product of hyperbolic tangents is the only continuously differentiable activation that performs equal or better than the ReLU activation, without any hyperparameter changes.
-Furthermore, using a product of hyperbolic tangent representations seems to greatly outperform the single hyperbolic tangent.
+In (a), the Tanh (HR) significantly outperforms a the baseline Tanh.
 In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown.
-Notably, the decrease in performance when using a single hyperbolic tangent with a twice as large representation ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable.
-We perform all experiments on 8 common, non-exploration driven Atari games.
-We normalize using the ReLU baseline normalized scores and show in Appendix D.1 that human-normalized scores provide an even better score for Tanh (LaN), due to the strong LaN performance in one of the games where humans perform much worse.
-Reinforcement Learning
-In DQN, the action $a_{t}$ is chosen following an $\epsilon$-greedy policy.
-With probability $\epsilon$, a random action is selected, and with $(1-\epsilon)$, the action maximizing the Q-value is chosen.
-The target $Y_{t}$ is defined as:
-$Y_{t}=r_{t}+\gamma Q^{\prime}(z_{t+1},\underset{a\in\mathcal{A}}{\operatorname*{arg\,max}}Q(z_{t+1},a)),$ (3)
-where $Q^{\prime}(z,a)$ denotes the target Q-network, an auxiliary network that stabilizes the learning by providing a stable target for $Q(z,a)$.
-The parameters of $Q^{\prime}$ are updated less frequently to enhance learning stability.
-The loss function for training the network is:
-$\mathcal{L}_{Q}=\big{|}Y_{t}-Q(z_{t},a)\big{|}^{2}.$ (4)
-Proximal Policy Optimization (PPO) operates on a different principle, utilizing policy gradient methods for policy improvement.
-PPO seeks to update the policy by maximizing an objective function while preventing large deviations from the previous policy through a clipping mechanism in the objective’s estimator.
-The clipped policy gradient loss $L^{CLIP}$ is defined as:
-$L^{CLIP}(\theta)=\mathbb{E}\left[\min(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right],$ (5)
-where $r_{t}(\theta)$ represents the ratio of the probabilities under the new policy versus the old policy, and $\hat{A}_{t}$ is the advantage estimate at timestep $t$.
-This clipped surrogate objective ensures gradual and stable policy updates.
-4.1 Atari - DQN
-(a) Dead Neurons
-(b) Effective Rank
-Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames).
-Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior when training RL.
-Furthermore, a product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
-Figure 7: Additional ablations run on DQN in the Atari domain. Tanh (+LaN) represents an addition rather than an elementwise product with a LaN, and Tanh (2xLaN) uses two Latent assistance Networks.
-The influence of a LaN on downstream performance can be seen in Fig.
-5.
-A significant improvement over the standard hyperbolic tangent baseline is seen, as well as an important improvement over the strong ReLU baseline.
+Notably, the decrease in performance when using a conventional representation with twice as much neurons ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable.
+Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
+Figure 8: Additional ablations run on DQN in the Atari domain. Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
 Examining further ablations in Fig.
 8 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance.
-Furthermore, taking a product of 3 hyperbolic tangents (2$\times$×LaN) also appears to enhance performance, though the improvement is not significantly different from using a single LaN.
-This could be the result of increased contracting behavior in the early stage of training due to subsequent multiplication of activations whose absolute values are $<1$.
-Further experiments combining the ReLU activation with a LaN can be found in Appendix C.3.
-Mitigating dying Hyperbolic Tangents
-As mentioned in Section 3, the number of dead neurons is defined as the amount of neurons that display the same output for any given observation $s_{t}$.
-The amount of dead neurons over time when training on Atari can be seen in  6(a).
-For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ are dead, while for the sigmoid and hyperbolic tangent activation this number is around 40%.
-When using a hyperbolic tangent with a LaN, a reduction in dead neurons as compared to using a single hyperbolic tangent or sigmoid can be observed.
-We credit this to the inherent ability of a product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 3.
-This reduction of long-term dead hyperbolic tangents when using the LaN is remarkable as it is also coinciding with a stronger divergence of values in the initial stages of training (see Fig.
-4).
-Increasing Effective Rank
-We calculate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig.
-6(b).
-As observed by Gulcehre et al.
-(2022), a representation activated by a hyperbolic tangent or, residing in the same category, a sigmoid, already has a relatively high effective rank compared to a representation activated by a ReLU.
-Furthermore, as can additionally be seen in supervised learning (see Fig.
-2), a LaN notably improves the effective rank of the hyperbolic tangent, which can be seen as a measure of network ‘expressivity’.
-We summarize our results qualitatively in Table 2.
-Table 2: High-level summary of the strengths of a LaN.
-Activation function Performance RL Few dead neurons Effective rank
-Tanh (LaN) ++ ++ ++
+Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product.
+We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$.
+For human-normalized scores, we refer the reader to Appendix D.1.
+Evaluating Hadamard representations on PPO
+To evaluate results on a policy-based algorithm, additional experiments are run using the PPO algorithm (Schulman et al., 2017).
+For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2).
+The results can be found in Fig.
+9.
+Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and the best performance.
+The Tanh (HR) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations can be relatively unreliable across algorithms.
+(a) Normalized Score
+(b) Effective Rank
+Figure 9: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
+Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations.
+As also indicated by Teney et al.
+(2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood.
+A qualitative overview of our results can be found in Table 3.
+Table 3: High-level summary of the strengths of a Hadamard representation.
+Activation function Performance Few dead neurons Effective rank
+Tanh (HR) ++ ++ ++
 Tanh - - +
@@
 Sigmoid - - -
-4.2 Atari - PPO
-Additional experiments are run using the PPO algorithm (Schulman et al., 2017).
-For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2).
-The results can be found in Fig.
-8.
-Similarly to DQN, the Tanh (LaN) exhibits the highest effective rank and the best performance.
-The Tanh (LaN) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations are relatively unreliable.
-(a) Normalized Score
-(b) Effective Rank
-Figure 8: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
+Increasing Representation Dimensionality
+A Hadamard representation increases the amount of trainable parameters in the encoder without increasing the dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
+This is an important feature as simply increasing the dimension of the representation to $\mathbb{R}^{1024}$ is not often effective (see Fig.
+7(b)).
+Even though the fraction of dying neurons decreases and the effective rank increases when using $z_{t}\in\mathbb{R}^{1024}$, it seems that the ’relative’ effective rank, defined as the ratio $\frac{rank}{dim(z_{t})}$, actually goes down (See Appendix.
+C.2).
+We therefore hypothesize that reducing dying neurons and increasing a representation’s effective rank is only superficially solved by expanding representation dimensionality.
 5 Related Work
@@
 Gulcehre et al.
-(2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank and do not suffer strongly from rank decay as training continues.
+(2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues.
+Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022).
 Other work by Lyle et al.
@@
 (2023) further studied plasticity injection for long-term training.
-In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Arnob et al., 2021).
+In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024).
 This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
-Network Architecture
+Network Architecture in RL
 The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject.
@@
 (2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997).
-Using these ideas, the Resnet was invented (He et al., 2016).
-More specific to RL, Henderson et al.
+Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018).
+Further work by Henderson et al.
 (2018) showed differences in RL performances over different network architectures and nonlinear activations.
@@
 6 Conclusions and Discussion
-We have proposed a novel architecture called the Latent assistance Network (LaN) that augments a networks’ final hidden layer by taking the element-wise product with a parallel, independent hidden layer.
-We discussed and empirically showed that this decreases the occurrence of dead neurons in the representation $z_{t}$, increases network expressiveness, and notably improves performance compared to a single continuously differentiable activation such as the hyperbolic tangent.
-Our results also show an important increase in performance over the dominant ReLU activation.
-This paper therefore remains, to the best of our knowledge, the first to show the possibility of elevating a continuously differentiable activation to and beyond the level of the ReLU.
-Additionally, it should be recognized that the employed DQN baseline was initially optimized specifically for the ReLU activation function.
+This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem.
+Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer.
+We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness.
+This approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
 As investigated by Kumar et al.
-(2021), a compressed representation activated by a hyperbolic tangent has a much higher effective rank than the ReLU and could therefore yield stronger network expressiveness.
-We therefore remain hopeful that the hyperbolic tangent can potentially replace the ReLU in a variety of architectures.
-Making strong claims in this field however remains challenging, as the success of pruning strategies show that network expressiveness might already be in abundance in some settings (Frankle & Carbin, 2019; Obando-Ceron et al., 2024).
-Future work could focus on further identifying the intricacies of this architecture in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in supervised learning.
-Also, we believe that an implementation of this architecture in a continual learning setting with similar evaluations as the work in Abbas et al.
+(2021), a compressed representation activated by a hyperbolic tangent attains a much higher effective rank than when activated by a ReLU, and will therefore yield stronger network expressiveness.
+In light of this, the hyperbolic tangent can potentially replace the ReLU in a variety of architectures where expressiveness is required without using deeper networks.
+Future work could focus on further identifying the intricacies of hyperbolic tangents in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in a supervised learning setting (see Fig.
+2).
+Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al.
 (2023) could be promising.
 7 Limitations
-Our work has focused on the encoders’ final hidden layer in reinforcement learning.
-As literature suggests that convolutional layers have an even stronger preference for ReLU activations (Jarrett et al., 2009), we have not conducted experiments testing a LaN on the convolutional activations of our encoders.
-Due to the heavy computational requirements of the Atari domain, limitations of our work include the use of a moderate amount of 5 seeds per environment, as well as the lack of a hyperparameter search for using the ‘Tanh (LaN)’ setting in our experiment section.
-A hyperparameter search could further strengthen our results, as the strong baseline is specifically tuned for the ReLU activation.
-Furthermore, as we perform experiments on DQN and PPO, integration of our Latent assistance Network into more sophisticated algorithms such as Rainbow (Hessel et al., 2018) would be interesting.
+Our research has focused on the encoders’ final hidden layer in reinforcement learning.
+Both literature and preliminary experiments suggest that convolutional layers have a very strong preference for rectified activations (Jarrett et al., 2009), possibly due to a low-frequency bias (Ramasinghe & Lucey, 2022; Teney et al., 2024).
+We have therefore not conducted experiments integrating Hadamard representations in the convolutional section of our encoder architectures.
+In addition, a hyperparameter search could further strengthen our results, since the baseline is specifically tuned for the ReLU activation.
+Furthermore, as we perform experiments on DQN and PPO, integration of a Hadamard representation into more sophisticated algorithms and architectures such as Rainbow (Hessel et al., 2018) and Impala (Espeholt et al., 2018) would be interesting.
+Reproducibility Statement
+Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1.
+The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022).
+Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
 Appendix A Implementation Details
@@
 Optimizer $\epsilon$ $10^{-5}$ Adam Epsilon
+Hadamard Implementation
+Constructing a Hadamard representation is a straightforward process that only requires an additional, parallel linear layer.
+Starting from the flattened convolutional features in DQN, the Pytorch pseudocode is defined as follows:
+flattened_features = conv_features.flatten(1)
+representation1 = nn.Tanh(linear1(flattened_features))
+representation2 = nn.Tanh(linear2(flattened_features))
+hadamard = representation1 * representation2
+Q-values = linear_q(hadamard)
+Reinforcement Learning
+In DQN, the action $a_{t}$ is chosen following an $\epsilon$-greedy policy.
+With probability $\epsilon$, a random action is selected, and with $(1-\epsilon)$, the action maximizing the Q-value is chosen.
+The target $Y_{t}$ is defined as:
+$Y_{t}=r_{t}+\gamma Q^{\prime}(z_{t+1},\underset{a\in\mathcal{A}}{\operatorname*{arg\,max}}Q(z_{t+1},a)),$ (3)
+where $Q^{\prime}(z,a)$ denotes the target Q-network, an auxiliary network that stabilizes the learning by providing a stable target for $Q(z,a)$.
+The parameters of $Q^{\prime}$ are updated less frequently to enhance learning stability.
+The loss function for training the network is:
+$\mathcal{L}_{Q}=\big{|}Y_{t}-Q(z_{t},a)\big{|}^{2}.$ (4)
+Proximal Policy Optimization (PPO) operates on a different principle, utilizing policy gradient methods for policy improvement.
+PPO seeks to update the policy by maximizing an objective function while preventing large deviations from the previous policy through a clipping mechanism in the objective’s estimator.
+The clipped policy gradient loss $L^{CLIP}$ is defined as:
+$L^{CLIP}(\theta)=\mathbb{E}\left[\min(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right],$ (5)
+where $r_{t}(\theta)$ represents the ratio of the probabilities under the new policy versus the old policy, and $\hat{A}_{t}$ is the advantage estimate at timestep $t$.
+This clipped surrogate objective ensures gradual and stable policy updates.
 A.2 PPO Architecture
-Figure 9: A visualisation of the Latent assistance Network (LaN) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment.
-A parallel linear layer providing an independent representation is integrated, where the element-wise product of the two parallel representations represents the final representation $z_{t}$.
+Figure 10: A visualisation of the Hadamard representation (HR) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment.
+A parallel linear layer providing an independent representation is integrated, where the Hadamard product of the two parallel representations represents the final representation $z_{t}$.
 Appendix B Kernel Density Estimations
-As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without a LaN are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons.
+As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without an HR are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons.
 We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig.
-10(a).
-(a) Tanh (LaN) - $5\cdot 10^{6}$ iterations
-(b) Tanh (LaN) - $5\cdot 10^{6}$ iterations
-(c) Tanh - $5\cdot 10^{6}$ iterations
-Figure 10: Kernel Density Estimations (KDE) over a subset of 16 neurons in the representations $z_{t}^{enc}$ and $z^{*}_{t}$ in (a), the resulting element-wise product $z_{t}$ in (b) and the representation $z_{t}$ when training without a LaN (c).
+11(a).
+(a) Tanh (HR) - $5\cdot 10^{6}$ iterations
+(b) Final Tanh (HR) - $5\cdot 10^{6}$ iterations
+(c) Tanh (no HR) - $5\cdot 10^{6}$ iterations
+Figure 11: Kernel Density Estimations (KDE) over a subset of 16 neurons in the representations $z_{t}^{enc}$ and $z^{*}_{t}$ in (a), the resulting Hadamard product $z_{t}$ in (b) and the representation $z_{t}$ when training without an HR (c).
 These representations are obtained after training DQN in the ’Breakout’ environment.
-A closer look at neurons 3, 8 and 9 show that when one of the representations saturates, the other is able to compensate, leading to a non-dead neuron in their product $z_{t}$.
+Red outlines represent dead (collapsed) neurons.
+In (a), a closer look at neurons 3, 8 and 9 shows that when one of the representations saturates, the other is able to compensate, leading to a non-dead neuron in their product $z_{t}$ in (b).
 (a) Tanh - $10^{7}$ iterations
-(b) Tanh (LaN) - $10^{7}$ iterations
+(b) Tanh (HR) - $10^{7}$ iterations
 (c) Sigmoid - $10^{7}$ iterations
-(d) Sigmoid (LaN) - $10^{7}$ iterations
+(d) Sigmoid (HR) - $10^{7}$ iterations
 Figure 12: Kernel Density Estimations of the final representation $z_{t}$ after training DQN for $10^{7}$ iterations in the Breakout environment.
-A hyperbolic tangent using LaN allows the representation to avoid strong saturation, keeping sufficient kernel density in the central sections of the hyperbolic tangent.
-As a sigmoid can saturate into zero, using a LaN remains less effective into preventing saturation.
+A hyperbolic tangent Hadamard representation allows the representation to avoid strong saturation, keeping sufficient kernel density in the central sections of the hyperbolic tangent.
+As a sigmoid can saturate into zero, using a Hadamard representation remains less effective for preventing saturation, as any zero will lead to a Hadamard product of zero.
 B.1 KDE calculation
@@
 where $\omega$ represents the predetermined threshold. In practice, after analyzing the individual neuron KDE’s, using an $\omega$ of 20 provides a strong approximation of actual dead neurons.
+B.2 Effective Rank calculation
+In line with Kumar et al.
+(2021), the effective rank of a feature matrix for a threshold $\delta$ ($\delta=0.01$), denoted as $srank_{\delta}(\Phi)$, is given by $srank_{\delta}(\Phi)=\min\left\{k:\frac{\sum_{i=1}^{k}\sigma_{i}(\Phi)}{\sum_{i=1}^{d}\sigma_{i}(\Phi)}\geq 1-\delta\right\}$, where $\{\sigma_{i}(\Phi)\}$ are the singular values of $\Phi$ in decreasing order, i.e., $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$.
+Intuitively, the effective rank of a feature matrix represents the number of “effective” unique components that form the basis for linearly approximating the resulting Q-values.
+The calculation in Python is done as follows:
+def compute_rank_from_features(feature_matrix, rank_delta=0.01):
+    sing_values = np.linalg.svd(feature_matrix, compute_uv=False)
+    cumsum = np.cumsum(sing_values)
+    nuclear_norm = np.sum(sing_values)
+    approximate_rank_threshold = 1.0 - rank_delta
+    threshold_crossed = (cumsum >= approximate_rank_threshold * nuclear_norm)
+    effective_rank = sing_values.shape[0] - np.sum(threshold_crossed) + 1
+    return effective_rank
 Appendix C Additional Experiments
 C.1 Shallow and Deep Function Approximation
-To further showcase the effect of activations on complex function approximation, we compare the single hidden layer Tanh (LaN) network from Fig.
+To further showcase the effect of activations on complex function approximation, we compare the single hidden layer Tanh (HR) network from Fig.
 2 with a deep ReLU and Tanh network containing three hidden layers each.
 The comparison with shallow networks can be found in Fig.
-12(a) and a comparison with deep networks can be found in Fig.
-12(b).
+13(a) and a comparison with deep networks can be found in Fig.
+13(b).
 (a) Comparison of shallow networks for a nonlinear regression task.
-The Tanh and ReLU networks have a single hidden layer of 200 neurons, while the Tanh (LaN) has a single hidden layer of 100 neurons but two preceding linear layers.
-The Tanh and ReLU networks have 601 parameters, while the Tanh (LaN) network has 501 parameters.
+The Tanh and ReLU networks have a single hidden layer of 200 neurons, while the Tanh (HR) has a single hidden layer of 100 neurons but two preceding linear layers.
+The Tanh and ReLU networks have 601 parameters, while the Tanh (HR) network has 501 parameters.
 As found by Gulcehre et al.
 (2022), a shallow network activated by ReLU has a lower effective rank and consequently reduced network expressivity as compared to a Tanh activated network.
-Using a LaN, we achieve better function approximation while using less parameters.
+Using a Hadamard representation, we achieve better function approximation while using less parameters.
 (b) Comparison of two deep networks and one shallow network for the same nonlinear regression task.
-The Tanh and ReLU networks have 3 hidden layers of 200 neurons each, while the Tanh (LaN) network remains shallow.
+The Tanh and ReLU networks have 3 hidden layers of 200 neurons each, while the Tanh (HR) network remains shallow.
 In line with common observations in deep learning, the ReLU activation thrives in deeper networks, in contrast to the Tanh activation.
-Interestingly, the shallow Tanh (LaN) network still achieves better function approximation with only 0.6% of the deeper networks’ parameters (81001 vs 501).
+Interestingly, the shallow Tanh (HR) network still achieves better function approximation with only 0.6% of the deeper networks’ parameters (81001 vs 501).
 No hyperparameter tuning or architecture search has been applied.
-Additional tests using deep Tanh (LaN) networks gave similar function approximation as compared to the shallow Tanh (LaN) network.
-C.2 Validation of dying neuron probability derivations
-As discussed in Section 3, the effect of adding a Latent assistance Network (LaN) to the original representation strongly depends on the activation function.
+Additional tests using deep Tanh (HR) networks gave similar function approximation as compared to the shallow Tanh (HR) network.
+C.2 Increasing Representation Parameters
+(a) Performance
+(b) Dead Neurons
+(c) Effective Rank
+Figure 14: Comparison of a normal hyperbolic tangent (Tanh), a hyperbolic tangent with a higher representation dimension $z_{t}\in\mathbb{R}^{512\rightarrow 1024}$ and a Hadamard representation using hyperbolic tangents.
+Comparisons are done on Performance (a), the fraction of dead neurons (b) and the effective rank of the representation $z_{t}$ (c).
+Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it can negatively impact actual convergence.
+C.3 Validating dying neuron probability derivations
+As discussed in Section 3, the effect of using a Hadamard representation strongly depends on the activation function.
 These derivations are empirically validated by the results in Fig.
-13.
+15.
 In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1.
-This could explain the very slight increase in dead neurons when using a LaN with a sigmoid.
-(a) Sigmoid with and without LaN
-(b) Tanh with and without LaN
-(c) ReLU with and without LaN
-Figure 13: By evaluating the effect of a LaN on dying neurons through the lens of probability theory, we predicted that only the hyperbolic tangent benefits in this metric.
+This could explain the very slight increase in dead neurons when using an HR with activations.
+(a) Sigmoid with and without HR
+(b) Tanh with and without HR
+(c) ReLU with and without HR
+Figure 15: By evaluating the effect of an HR on dying neurons through the lens of probability theory, we predicted that only the hyperbolic tangent benefits in this metric.
 Specifically, only a hyperbolic tangent was speculated to have a decrease in dying neurons.
-Using a LaN on a sigmoid activation would have no notable difference, and for the ReLU activation we expected an increase in dead neurons.
-This empirically validates our theory in Section 3.
-C.3 Combining ReLU with a LaN
-Additional Atari experiments are provided comparing a ReLU activation with and without a LaN. The results can be seen in fig. 14.
+Using an HR with sigmoid activations would have no notable difference, and for an HR with ReLU activations an increase in dead neurons was expected.
+This empirically validates our hypotheses in Section 3.
+C.4 ReLU activated Hadamard representation
+Additional Atari experiments are provided comparing a ReLU activation with and without an HR. The normalized scores, dying neurons and the effective rank during training can be seen in fig. 16.
 (a) Normalized Score
@@
 (c) Effective Rank
-Figure 14: As a Rectified Linear Unit creates sparse representations, it does not benefit from using a LaN, since the final representation will consist of the element-wise product between two sparse representations.
+Figure 16: As a Rectified Linear Unit creates sparse representations, it does not benefit from using an HR, since the final representation will consist of the Hadamard product between two sparse representations.
 Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
@@
 Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig.
-15.
+17.
 Due to taking a subset of the full atari domain, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation.
 For a more realistic comparison of the methods, we therefore decided to use baseline-normalized scores in the main paper.
-Figure 15: Human-Normalized performance (in multiples) with the standard deviation over the means in the Atari domain for 10M iterations (40M Frames)
-D.2 Individual Environments
+Figure 17: Human-Normalized performance (in multiples) with the standard deviation over the means in the Atari domain for 10M iterations (40M Frames).
+D.2 Individual Environment Scores
 Figure 18: DQN Performance comparison on the individual Atari Environments.
@@
 Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations.
-To avoid cherry-picking, no other Atari games were tested.
+No other games were tested.
 Figure 19: PPO Performance comparison on the individual Atari Environments.
@@
 Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations.
-To avoid cherry-picking, no other Atari games were tested.
+No other games were tested.

### v3

@@
 The most commonly used activation functions can be classed into the category of continuously differentiable (e.g.
-tanh) and linear-unit functions (e.g.
+tanh) and piece-wise linear functions (e.g.
 ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g.
 measured by the number of dead neurons and the effective rank).
-In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions.
+In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to piece-wise linear functions.
 We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s.
 To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation.
-Using deep Q-networks and proximal policy optimization in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
+Using deep Q-networks, proximal policy optimization and parallelized Q-networks in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
 1 Introduction
@@
 As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
-In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018).
+In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023).
+However, even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018).
 Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig.
@@
 Based on these insights, we mitigate said effects by augmenting the original architecture.
-Specifically, we focus our research on the activations of the compressed representation of an encoder, defined as the final hidden layer, and provide an alternative to the conventional parameterization of this representation.
+Specifically, we focus our research on the activations of an encoder, defined as the hidden layer, and provide an alternative to the conventional parameterization of this representation.
 Our contributions can be summarized as follows:
-  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem.
+  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, but have a more profound effect on performance.
   • A Hadamard representation (HR) is proposed, defining a latent representation as the Hadamard product of two separate, individually parameterized activation vectors.
-  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, a Hadamard representation yields significant performance gains over a standard hyperbolic tangent in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
+  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, using Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
 2 Preliminaries
@@
 3 Augmenting Hyperbolic Tangents
-Continuously differentiable activations such as the hyperbolic tangent ($tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the $ReLU$ or its Linear-Unit descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity).
+Continuously differentiable activations such as the hyperbolic tangent ($\tanh$) and the sigmoid ($\sigma$) activations are fundamentally different than the ReLU or its piece-wise linear descendants, which are non-symmetric and have a large part of the input space mapped to zero (leading to sparsity).
 The hyperbolic tangent and the sigmoid output values in the ranges $[-1,1]$ and $[0,1]$, respectively.
-These functions are defined as:
-$tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}},\qquad\sigma(x)=\frac{1}{1+e^{-x}}$ (1)
+These functions are defined as $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ and $\sigma(x)=\frac{1}{1+e^{-x}}$.
 Both functions have the advantage of being differentiable everywhere, as well as being bounded.
@@
 Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior.
-When using a hyperbolic tangent in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, can be classified as dying if:
-$|\alpha_{i}|\approx 1\;\forall\;s_{t}\in B$ (2)
-Where $s_{t}$ is an observation in buffer $B$.
-Note that an absolute value is used, as a hyperbolic tangent can either saturate towards -1 or 1.
-Given that the hyperbolic tangent is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) .
+When using any activation function in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, is saturated or dying if:
+$\alpha_{i}\approx\Omega\;\forall\;s_{t}\in B$ (1)
+Where $\Omega$ represents the saturation value and $s_{t}$ is an observation in buffer $B$.
+In practice, a mini-batch of observations is evaluated instead of the whole dataset in $B$.
+For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) .
 To approximate the condition given in Eq.
-2, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer.
+1, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer.
 In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken.
@@
 In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$.
-This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015).
+This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015), or as an augmented version of the Gated Linear Unit (GLU) (Dauphin et al., 2017).
 The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$.
@@
 Vertical and horizontal bars represent convolutional and linear layers, respectively.
-A parallel linear layer providing independently parameterized hidden layer activations is integrated, where the Hadamard product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
+A parallel hidden layer providing independently parameterized activations is integrated, where the Hadamard product of the two hidden layers represents the final latent representation $z_{t}$ used for downstream learning.
 Preventing Dying Neurons
 Our key hypothesis is that the Hadamard representation can prevent saturation, hence alleviating vanishing gradients and dying neurons (See Eq.
-2).
+1).
 To support our hypothesis, we investigate the derivative of a product of two functions.
@@
 For a product of two hyperbolic tangent functions, the derivative is defined as:
-$h^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
-where $\operatorname{sech}^{2}$ is the derivative of a hyperbolic tangent.
-When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values.
-This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ is still able to maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients.
+$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
+In this context, $\operatorname{sech}^{2}$ is the derivative of the hyperbolic tangent function.
+Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occours.
+Thus, when $g(x)$ saturates, $h(x)$ still keeps a non-trivial gradient in the product, providing a mechanism to avoid vanishing gradients.
 We visualize the kernel densities during training with a Hadamard representation in Fig.
@@
 Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$.
-Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron’s activation function.
+Lastly, we make an independence assumption between two individual neurons.
+Under these assumptions, we show that interpreting a neuron as the product of two individual neurons can change saturation probabilities depending on the neuron’s activation function.
 (a) Tanh (HR) - $10^{6}$ iterations
@@
 4 Experiments
-The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017).
-In all experiments, we build on the implementations from cleanrl  (Huang et al., 2022).
-For more details on the PPO and DQN implementation, we refer the reader to Appendix A.1.
-For both DQN and PPO, we define the representation $z_{t}\in\mathbb{R}^{512}$ as the last hidden layer of the network, which serves as the compressed representation of the original pixel observation $s_{t}$ (see Fig.
-4).
-We perform all experiments on 8 common, non-exploration driven Atari games.
+The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017) and the parallelized Q-network (PQN) (Gallici et al., 2024).
+For more details on the PPO, DQN and PQN implementation, we refer the reader to Appendix A.1.
+In line with Gulcehre et al.
+(2022), for our qualitative experiments, a Hadamard Representation is only applied to the last hidden layer $z_{t}\in\mathbb{R}^{512}$.
+We perform qualitative experiments on 8 common, non-exploration driven Atari games.
 We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance.
@@
 The performance is aggregated over 8 environments and normalized using the ReLU baseline scores.
+Finally, the PQN experiments are conducted with an HR on all hidden layers, showing the Median Human-Normalized scores on 51 Atari games.
 (a) Dead Neurons
@@
 As mentioned in Section 3 and defined in Eq.
-2, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$.
+1, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$.
 The amount of dead neurons over time when training on Atari can be seen in Fig.
@@
 ReLU 0.62 0.73 +18%
-For more dying neuron graphs related to the ReLU and the Sigmoid activations, we refer the reader to Appendix C.3.
+For more activation-specific dying neuron graphs, we refer the reader to Appendix C.3.
 Increasing Effective Rank
@@
 Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected.
-For more information on the effect of a Hadamard representation on a ReLU, we refer the reader to Appendix C.4.
 Analyzing Performance in Atari
@@
 7.
-Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement of nearly $60\%$ over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline.
-Note that no hyperparameter or architectural tuning has been employed in our experiments.
+Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline.
+Furthermore, a comparison is made with the novel Rational (Delfosse et al., 2024) activation function as the activation in the final hidden layer.
+Fig.
+7(a) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
 (a) Normalized Score
@@
 Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames).
-In (a), the Tanh (HR) significantly outperforms a the baseline Tanh.
-In (b), several ablations including layer normalization (Ba et al., 2016) on hyperbolic tangent activations are shown.
-Notably, the decrease in performance when using a conventional representation with twice as much neurons ($z_{t}\in\mathbb{R}^{1024}$) shows that more parameters are not always desirable.
-Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
+In (a), the Tanh (HR) significantly outperforms the baseline Tanh.
+The recently proposed Rational Activation (Delfosse et al., 2024) seems to have comparable performance to the ReLU in the 40M frame setting.
+In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced $\alpha=5e^{-5}$ on hyperbolic tangent activations are shown.
+Note that no hyperparameter changes are done for the Hadamard Representation.
 Figure 8: Additional ablations run on DQN in the Atari domain. Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
 Examining further ablations in Fig.
-8 shows that using a sophisticated Linear-Unit function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance.
+8 shows that using a sophisticated piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance.
 Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product.
 We hypothesize that this is the result of increased contracting behavior in the early stage of training due to increasing multiplication of hyperbolic tangent activations whose absolute values are $<1$.
-For human-normalized scores, we refer the reader to Appendix D.1.
+Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
 Evaluating Hadamard representations on PPO
@@
 Figure 9: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
+Evaluating Hadamard representations on PQN
+To more reliably test the performance gains one can achieve using Hadamard Representations, we evaluate their effects on the Parallelized Q-Network (PQN) in the full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
+Figure 10: Median Human-Normalized scores on PQN in 51 Atari games for 5 seeds over 40M frames.
+Labels represent MLP / CNN activations.
+Employing a Hadamard Representation in the Convolutional layers as well as the MLP layer provides a significant performance improvement over the recent PQN baseline, as well as a CReLU (Abbas et al., 2023) implementation of PQN.
+PQN has recently emerged as a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024).
+We run tests implementing the Hadamard Representation in all hidden layers.
+The median human-normalized scores after training for 40M frames are shown in Fig.
+10.
+We can clearly see a similar trend for the 51-game Atari suite, where using a Hadamard representation for all layers significantly increases the median performance as compared to the baseline PQN and CReLU (Abbas et al., 2023)
+Hyperbolic Tangents turn Weights into Biases
 Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations.
@@
 (2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood.
-A qualitative overview of our results can be found in Table 3.
-Table 3: High-level summary of the strengths of a Hadamard representation.
-Activation function Performance Few dead neurons Effective rank
-Tanh (HR) ++ ++ ++
-Tanh - - +
-ReLU +
-Sigmoid - - -
-Increasing Representation Dimensionality
-A Hadamard representation increases the amount of trainable parameters in the encoder without increasing the dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
-This is an important feature as simply increasing the dimension of the representation to $\mathbb{R}^{1024}$ is not often effective (see Fig.
-7(b)).
-Even though the fraction of dying neurons decreases and the effective rank increases when using $z_{t}\in\mathbb{R}^{1024}$, it seems that the ’relative’ effective rank, defined as the ratio $\frac{rank}{dim(z_{t})}$, actually goes down (See Appendix.
-C.2).
-We therefore hypothesize that reducing dying neurons and increasing a representation’s effective rank is only superficially solved by expanding representation dimensionality.
+However, in the case of dying neurons, we aim to shed some light on this phenomenon.
+When dying neurons occur in ReLU-activated layers, it basically prunes these neurons and the associated weights to the next layer.
+However, in hyperbolic tangent activated layers, dying neurons lead to an unintended phenomenon where weights associated with dead neurons effectively become biases.
+Theorem 4.1.
+When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons, inducing a hidden bias.
+Proof.
+Take a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\bm{w}^{j}_{\alpha_{i}}$ in layer $z^{j}$. The influence of these neurons on the next layer $z^{j+1}$ is calculated as:
+$z^{j+1}=\sum_{i}\alpha^{j}_{i}\bm{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
+If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0~{}\forall~{}s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
+$\sum_{i}0\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=0\quad\forall s\in\mathcal{S}.$ (3)
+Where $\bm{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons.
+However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\}\forall s\in\mathcal{S}$), the output is
+$\sum_{i}\{1,-1\}\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1}\quad\forall s\in\mathcal{S}.$ (4)
+As a result, the weight vector $\bm{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the resulting hidden layer.
+Note that the bias $B^{j+1}_{*}$ is constant for any input observation.
+∎
+This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ next hidden layer or Q-values, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning.
 5 Related Work
@@
 Nikishin et al.
-(2023) further studied plasticity injection for long-term training.
+(2023) further studied plasticity injection for long-term training and Delfosse et al.
+(2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity.
+Concurrent work by Dohare et al.
+(2024) used continual back-propagation to further alleviate plasticity loss.
 In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024).
@@
 We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness.
-This approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
-As investigated by Kumar et al.
-(2021), a compressed representation activated by a hyperbolic tangent attains a much higher effective rank than when activated by a ReLU, and will therefore yield stronger network expressiveness.
-In light of this, the hyperbolic tangent can potentially replace the ReLU in a variety of architectures where expressiveness is required without using deeper networks.
-Future work could focus on further identifying the intricacies of hyperbolic tangents in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in a supervised learning setting (see Fig.
+In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
+Finally, we have given insights into the ability of the ReLU activation to be less affected the symptoms of dying neurons, as opposed to non-zero saturating activations such as the hyperbolic tangent.
+Future work could focus on further identifying the intricacies of the effects that different activation functions have on the resulting representation, in an attempt to push the potential of non piece-wise linear activations in reinforcement learning or even in a supervised learning setting (see Fig.
 2).
 Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al.
+(2023) or in Delfosse et al.
 (2024) could be promising.
 7 Limitations
-Our research has focused on the encoders’ final hidden layer in reinforcement learning.
-Both literature and preliminary experiments suggest that convolutional layers have a very strong preference for rectified activations (Jarrett et al., 2009), possibly due to a low-frequency bias (Ramasinghe & Lucey, 2022; Teney et al., 2024).
-We have therefore not conducted experiments integrating Hadamard representations in the convolutional section of our encoder architectures.
-In addition, a hyperparameter search could further strengthen our results, since the baseline is specifically tuned for the ReLU activation.
-Furthermore, as we perform experiments on DQN and PPO, integration of a Hadamard representation into more sophisticated algorithms and architectures such as Rainbow (Hessel et al., 2018) and Impala (Espeholt et al., 2018) would be interesting.
+Using a Hadamard representation will double the incoming weights connected to a hidden layer.
+However, in Fig.
+10, a comparison against the CReLU is shown, which equally doubles the network’s parameters.
+Furthermore, recent work shows that a scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024).
+Second, due to the vast amount of activation functions and the theoretical sound candidacy for the hyperbolic tangent, we have mostly focused our research on the combination of hyperbolic tangents.
+Further research into multiplicative representations could find even better activation combinations.
+A hyperparameter search could also give more insights into the strengths of the Hadamard Representation, since the baselines are specifically tuned for the ReLU activation.
+Finally, as this research performs experiments on DQN, PPO and PQN, integration of a Hadamard representation into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al.
+(2021) would be interesting.
 Reproducibility Statement
@@
 The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022).
+The PQN algorithm uses the implementation in the Github repository of Gallici et al.
+(2024).
 Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
@@
 Optimizer $\epsilon$ $10^{-5}$ Adam Epsilon
+Table 5: PQN Hyperparameters (Gallici et al., 2024)
+Hyperparameter Value Description
+Total Timesteps 10,000,000 Total timesteps for training
+Timesteps for Decay 10,000,000 Timesteps for decay functions (epsilon and lr)
+Number of Environments 128 Number of parallel environments
+Steps per Environment 32 Steps per environment in each update
+Number of Epochs 2 Number of epochs per update
+Number of Minibatches 32 Number of minibatches per epoch
+Starting Epsilon 1.0 Starting epsilon for exploration
+Final Epsilon 0.001 Final epsilon for exploration
+Epsilon Decay Ratio 0.1 Decay ratio for epsilon
+Epsilon for Test Policy 0.0 Epsilon for greedy test policy
+Learning Rate 0.00025 Learning rate
+Learning Rate Linear Decay True Use linear decay for learning rate
+Max Gradient Norm 10.0 Max gradient norm for clipping
+Discount Factor $(\gamma)$ 0.99 Discount factor for reward
+Lambda $(\lambda)$ 0.65 Lambda for generalized advantage estimation
+Episodic Life True Terminate episode when life is lost
+Reward Clipping True Clip rewards to range [-1, 1]
+Frame Skip 4 Number of frames to skip
+Max No-Ops on Reset 30 Max number of no-ops on reset
+Test During Training True Run evaluation during training
+Number of Test Envs 8 Number of environments used for testing
 Hadamard Implementation
@@
 We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig.
-11(a).
+12(a).
 (a) Tanh (HR) - $5\cdot 10^{6}$ iterations
@@
 The comparison with shallow networks can be found in Fig.
-13(a) and a comparison with deep networks can be found in Fig.
-13(b).
+14(a) and a comparison with deep networks can be found in Fig.
+14(b).
 (a) Comparison of shallow networks for a nonlinear regression task.
@@
 (c) Effective Rank
+(d) Learning Rates $z_{t}\in\mathbb{R}^{1024}$
 Figure 15: Comparison of a normal hyperbolic tangent (Tanh), a hyperbolic tangent with a higher representation dimension $z_{t}\in\mathbb{R}^{512\rightarrow 1024}$ and a Hadamard representation using hyperbolic tangents.
-Comparisons are done on Performance (a), the fraction of dead neurons (b) and the effective rank of the representation $z_{t}$ (c).
-Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it can negatively impact actual convergence.
+Comparisons are done on performance (a), the fraction of dead neurons (b), the effective rank of the representation $z_{t}$ (c) and learning rates of the higher-dimensional latent.
+Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it often requires different hyperparameters, and can lead to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024).In (d), an ablation of learning rates shows that using a larger layer can sometimes prefer lower learning rates.
+However, it also shows that the improvement due to the Hadamard representation is likely not correlated with the parameter increase, as the Hadamard representation still significantly outperforms any of the 1024-dimensional latent state learning rate ablations.
 C.3 Validating dying neuron probability derivations
@@
 These derivations are empirically validated by the results in Fig.
-15.
+16.
 In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1.
 This could explain the very slight increase in dead neurons when using an HR with activations.
+Note that, since we use neuron independence assumptions in our theoretically calculated dying neuron probabilities, the empirical results differ in magnitude from the theoretical predictions.
 (a) Sigmoid with and without HR
@@
 Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
+C.5 Hyperbolic Tangent Convolutional Activation
+We additionally run tests in DQN on 8 Atari games. The baseline ReLU is used, except for the convolutional activations which are changed to a hyperbolic tangent.
+Figure 18: Baseline normalized performance in the Atari domain for 10M iterations (40M Frames) on DQN.
+It is evident that using non-Hadamard hyperbolic tangents in the convolutional layers can lead to only a fraction of the performance of the ReLU-activated baseline.
 Appendix D Atari
-D.1 Evaluation Details
+D.1 DQN Evaluation Details
 We normalize performance with respect to the ReLU baseline in cleanrl  Huang et al.
@@
 Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig.
-17.
-Due to taking a subset of the full atari domain, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation.
+19.
+Due to taking a subset of the Atari domain in DQN and PPO, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation.
 For a more realistic comparison of the methods, we therefore decided to use baseline-normalized scores in the main paper.
@@
 No other games were tested.
+Figure 22: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 23: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 24: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 25: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 26: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 27: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 28: PQN Performance comparison on the individual Atari Environments.
+Labels represent MLP / CNN activations.
+Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.

### v4

@@
 tanh) and piece-wise linear functions (e.g.
-ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g.
-measured by the number of dead neurons and the effective rank).
+ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning.
 In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to piece-wise linear functions.
-We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s.
-To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation.
-Using deep Q-networks, proximal policy optimization and parallelized Q-networks in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
+We show that the dying neuron problem in RL is not exclusive to ReLUs and actually leads to additional problems in the case of continuously differentiable activations such as tanh.
+To alleviate the dying neuron problem with these activations, we propose a Hadamard representation that unlocks the advantages of continuously differentiable activations.
+Using DQN, PPO and PQN in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
+Machine Learning, ICML
 1 Introduction
-Figure 1: Normalized performance training DQN in the Atari domain where the nonlinear activation function of the representation, defined as the final hidden layer, is either ReLU or tanh.
-A strong performance discrepancy due to a relatively small architectural change can be observed.
-Out of all activation functions, the Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016).
+Figure 1: Median Human-Normalized performance training PQN in the Atari domain where the activation function of the hidden layers is changed.
+In PQN, all activations are already layer-normalized.
+A massive performance discrepancy in performance can be observed when selecting different activation functions.
+Notably, in Atari, the application of a Hadamard representation with hyperbolic tangent leads to over 100% performance gains.
+The Hadamard representation is not suitable for the ReLU activation, as it amplifies dying neurons by taking the product of sparse activations.
+The Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016).
 The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
@@
 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019).
-As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
-In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023).
-However, even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018).
-Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig.
+As training progresses, the number of dying ReLUs tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
+In reinforcement learning (RL) (Sutton & Barto, 2018), the dying neuron phenomenon is much more prevalent than in supervised learning due to the use of non-stationary targets (Sokar et al., 2023).
+However, even though training results in a large number of dying ReLUs (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018).
+Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not favored in RL (see Fig.
 1).
@@
 Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
-We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity.
-Based on these insights, we mitigate said effects by augmenting the original architecture.
-Specifically, we focus our research on the activations of an encoder, defined as the hidden layer, and provide an alternative to the conventional parameterization of this representation.
-Our contributions can be summarized as follows:
-  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, but have a more profound effect on performance.
-  • A Hadamard representation (HR) is proposed, defining a latent representation as the Hadamard product of two separate, individually parameterized activation vectors.
-  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, using Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
-2 Preliminaries
 Figure 2: A regression of three shallow neural network architectures on a random complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$).
@@
 For experiments comparing deeper networks, we refer the reader to Appendix C.1.
-We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$.
-$\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor.
+This paper provides insights into the hyperbolic tangent’s suboptimality, revealing that RL gradients lead to bias-inducing dying neurons and under-utilization of the full network capacity.
+Based on these insights, we mitigate said effects by augmenting the original hidden layer architecture.
+Specifically, we provide an alternative to the conventional parameterization of these layers.
+Our contributions can be summarized as follows:
+  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem, and argue how they inherently have a more profound effect on performance.
+  • A Hadamard representation (HR) is proposed, defining a hidden layer as the Hadamard product of two separate, individually parameterized activation vectors.
+  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, Hadamard representations yields notable performance gains in multiple algorithms in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
+2 Related Work
+This section provides related work from the perspective of: (i) the loss of capacity through learning neural networks and (ii) the effect of different network architectures in RL.
+Network Capacity in RL.
+Liu et al.
+(2019) investigated the need for sparse representations in the continuous control domain.
+Gulcehre et al.
+(2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues.
+Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022).
+Other work by Lyle et al.
+(2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays.
+Further research by Nikishin et al.
+(2022) used network resets to counteract the primacy bias and (Sokar et al., 2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting.
+Nikishin et al.
+(2023) further studied plasticity injection for long-term training and Delfosse et al.
+(2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity.
+Concurrent work by Dohare et al.
+(2024) used continual back-propagation to further alleviate plasticity loss.
+In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024).
+This provides insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
+Network Architecture in RL
+The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject.
+Work by (Srivastava et al., 2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997).
+Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018).
+Further work by Henderson et al.
+(2018) showed differences in RL performances over different network architectures and nonlinear activations.
+Work by Abbas et al.
+(2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch.
+Finally, recent work by Grooten et al.
+(2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
+3 Preliminaries
+(a) Tanh - $10^{6}$ iterations
+(b) Tanh - $10^{7}$ iterations
+Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN (Mnih et al., 2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$.
+Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
+Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
+We consider an agent acting within its environment as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$.
+$\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{P}(\mathcal{S})$ is the environment’s transition function, $R:\mathcal{S}~{}\times~{}\mathcal{A}~{}\rightarrow~{}\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor.
 A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$.
@@
 The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
-Furthermore, we examine the setting where a high-dimensional state ($s_{t}\in\mathbb{R}^{v}$) is compressed into lower-dimensional activations $z_{t}\in\mathcal{Z}=\mathbb{R}^{w}$ where we call $\mathcal{Z}$ the representation space with $w\leq v$.
-This is done by means of a neural network encoding $e:\mathcal{S}\rightarrow\mathcal{Z}$ where $e$ represents the encoder.
 4 Augmenting Hyperbolic Tangents
@@
 However, both functions exhibit the vanishing gradient problem for saturating activations (Glorot & Bengio, 2010; Goodfellow et al., 2016).
+Figure 4: 10M iterations (40M frames) training DQN with a hyperbolic tangent activation in the Seaquest environment.
+The average contribution to the Q-values of the live and dead neurons in the final hidden layer is observed.
+If a neuron dies, it retains the same value for any input observation, but a multiplication of the nonzero saturation value with its outgoing weights implements a substantial ‘hidden’ bias on the Q-values.
 Dying Hyperbolic Tangents
-Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior.
+Although the literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior, particulary in the RL context with the moving target from the Bellman iterations.
 When using any activation function in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, is saturated or dying if:
-$\alpha_{i}\approx\Omega\;\forall\;s_{t}\in B$ (1)
+$\alpha_{i}\approx\Omega,\;\forall s_{t}\in B$ (1)
+Figure 5: A visualisation of the Hadamard representation.
+Horizontal bars represent weight vectors and $\boldsymbol{z}_{t}$ represents a hidden layer.
+Between each hidden layer, two parallel independently parameterized activation layers are formed, where the Hadamard product of the two activation layers represents the actual propagated hidden layer.
 Where $\Omega$ represents the saturation value and $s_{t}$ is an observation in buffer $B$.
 In practice, a mini-batch of observations is evaluated instead of the whole dataset in $B$.
-For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) .
+For the hyperbolic tangent, given that it is an asymptotic function near its saturation point, an approximate equality is considered ($|\alpha_{i}|\neq 1,\;\forall s_{t}\in B$) .
 To approximate the condition given in Eq.
@@
 3.
-(a) Tanh - $10^{6}$ iterations
-(b) Tanh - $10^{7}$ iterations
-Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN Mnih et al.
-(2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$.
-Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$.
-Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
+Massive KDE spikes at either 1 or -1 represent the absence of neuron variance over a batch of observations, meaning that the neuron has lost its value to propagate useful information.
 More details on the KDE calculation and dying neuron classification can be found in Appendix B.
+Hyperbolic Tangents turn Weights into Biases
+When dying neurons occur in ReLU-activated layers, it basically prunes these neurons and the associated weights to the next layer.
+However, in hyperbolic tangent activated layers, dying neurons lead to an unintended phenomenon where weights associated with dead neurons effectively become biases.
+Theorem 4.1.
+When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons multiplied by their corresponding weights, resulting in the hidden bias $B_{*}^{j+1}$.
+Proof.
+Let us consider a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\boldsymbol{w}^{j}_{\alpha_{i}}$ in the hidden layer $z^{j}$.
+The influence of these neurons on the next hidden layer $z^{j+1}$ is calculated as:
+$z^{j+1}=\sum_{i}\alpha^{j}_{i}\boldsymbol{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
+If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0,~{}\forall s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
+$\sum_{i}0\cdot\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}=0,\quad\forall s\in\mathcal{S}.$ (3)
+Where $\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons.
+However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\},\forall s\in\mathcal{S}$), the output is
+$\sum_{i}\{1,-1\}\cdot\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1},\quad\forall s\in\mathcal{S}.$ (4)
+As a result, the weight vector $\boldsymbol{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the next layer.
+Note that the bias $B^{j+1}_{*}$ is constant for any input observation.
+∎
+This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ hidden layers or output layer, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning.
+For example, in Atari’s Seaquest environment, we can see that this bias is substantial and nonzero (see Fig.
+4).
 Hadamard Representations (HR)
@@
 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed.
-In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer.
+In the conventional encoder setting, a network’s hidden layer can be defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer.
 In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$.
 This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015), or as an augmented version of the Gated Linear Unit (GLU) (Dauphin et al., 2017).
-The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$.
+The final representation is defined as a Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$.
 A visualization of the proposed architecture can be found in Fig.
-4.
-Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015).
-Vertical and horizontal bars represent convolutional and linear layers, respectively.
-A parallel hidden layer providing independently parameterized activations is integrated, where the Hadamard product of the two hidden layers represents the final latent representation $z_{t}$ used for downstream learning.
+5.
 Preventing Dying Neurons
@@
 To support our hypothesis, we investigate the derivative of a product of two functions.
-For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}g(x)h^{\prime}(x)}}$.
+For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g(x)h^{\prime}(x)}}$.
 In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
-$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}$
-$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
+$z^{\prime}(x)=\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
+$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
 If a neuron from $f(A_{1}x+B_{1})=0\hskip 2.84526pt\forall\hskip 2.84526ptx$, the gradient of the product becomes 0 while if a positive saturation is experienced i.e.
@@
 For a product of two hyperbolic tangent functions, the derivative is defined as:
-$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
+$z^{\prime}(x)$ $=\uline{\color[rgb]{0.40625,0.1953125,0.97265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.40625,0.1953125,0.97265625}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{1}\,\text{sech}^{2}(A_{1}x+B_{1})\,\tanh(A_{2}x+B_{2})}}$
+$\quad+\,\uline{\color[rgb]{0.296875,0.76171875,0.83203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{2}\,\text{sech}^{2}(A_{2}x+B_{2})\,\tanh(A_{1}x+B_{1})}}.$
 In this context, $\operatorname{sech}^{2}$ is the derivative of the hyperbolic tangent function.
-Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occours.
+Unlike the sigmoid, the hyperbolic tangent saturates to nonzero values, ensuring that if and only if both parts are saturated, product saturation occurs.
 Thus, when $g(x)$ saturates, $h(x)$ still keeps a non-trivial gradient in the product, providing a mechanism to avoid vanishing gradients.
 We visualize the kernel densities during training with a Hadamard representation in Fig.
-5.
+6.
 The individual representations before taking the Hadamard product can be found in Appendix B.
@@
 Under these assumptions, we show that interpreting a neuron as the product of two individual neurons can change saturation probabilities depending on the neuron’s activation function.
-(a) Tanh (HR) - $10^{6}$ iterations
-(b) Tanh (HR) - $10^{7}$ iterations
-Figure 5: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the Breakout environment using a Hadamard representation (HR) with hyperbolic tangents.
-The Hadamard representation tends to quickly utilize the full range of the hyperbolic tangent while also mitigating dying neurons.
 Hyperbolic Tangent:
@@
 This results in a probability of $\mathbf{p}\cdot\mathbf{p}=\mathbf{p}^{2}$.
-Taking a product of hyperbolic tangent activated neurons thus reduces the probability of neuron saturation from $\mathbf{p}$ to $\mathbf{p}^{2}$.
+Taking a product of hyperbolic tangent activated neurons thus reduces the neuron saturation probability from $\mathbf{p}$ to $\mathbf{p}^{2}$.
 Sigmoid:
@@
 Rectified Linear-Unit:
-In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$.
+In the case of a ReLU activation, the probability of a single neuron dying is taken as $\mathbf{p}$.
 As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$.
@@
 Table 1: Predicted dying neuron probabilities with and without a Hadamard representation.
-Activation Function Dying Neuron Probability Probability with HR Difference
-Hyperbolic Tangent $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
+Activation Prob. Prob. with HR $\Delta$
+Tanh $\mathbf{p}$ $\mathbf{p}^{2}$ $-(\mathbf{p}-\mathbf{p}^{2})$
 Sigmoid $\mathbf{p}$ $\mathbf{p}$ 0
@@
 5 Experiments
-The main experiments and discussion are conducted with the value-based DQN algorithm (Mnih et al., 2015) in Atari, and complemented by an evaluation on the policy-based proximal policy optimization algorithm (PPO) (Schulman et al., 2017) and the parallelized Q-network (PQN) (Gallici et al., 2024).
-For more details on the PPO, DQN and PQN implementation, we refer the reader to Appendix A.1.
-In line with Gulcehre et al.
-(2022), for our qualitative experiments, a Hadamard Representation is only applied to the last hidden layer $z_{t}\in\mathbb{R}^{512}$.
-We perform qualitative experiments on 8 common, non-exploration driven Atari games.
-We aim to show the effect of a Hadamard representation on the representation’s fraction of dying neurons, its effective rank and its downstream performance.
-For more information on the dying neuron and effective rank calculations, we refer the reader to Appendix.
-B.1 and Appendix.
-B.2, respectively.
-The performance is aggregated over 8 environments and normalized using the ReLU baseline scores.
-Finally, the PQN experiments are conducted with an HR on all hidden layers, showing the Median Human-Normalized scores on 51 Atari games.
-(a) Dead Neurons
-(b) Effective Rank
-Figure 6: The average fraction of dead neurons (a) and the average effective rank (Kumar et al., 2021) (b) of the representation $z_{t}$ when training DQN in the Atari domain for 10M iterations (40M frames).
-Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior during training.
+(a) Tanh (HR) - $10^{6}$ iterations
+(b) Tanh (HR) - $10^{7}$ iterations
+Figure 6: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the Breakout environment using a Hadamard representation (HR) with hyperbolic tangents.
+The Hadamard representation tends to quickly utilize the full range of the hyperbolic tangent while also mitigating dying neurons.
+We analyze the effect of a Hadamard representation on a hidden layer’s fraction of dying neurons, its effective rank and the downstream performance.
+In the qualitative analysis, we evaluate DQN (Mnih et al., 2015) and PPO (Schulman et al., 2017) on 8 common, non-exploration driven Atari environments for 40M frames.
+Finally, a larger performance analysis is done using the recent Parallelized Q-Network (PQN) (Gallici et al., 2024) on 51 Atari environments for 40M frames.
+(a) Dead Neurons (DQN)
+(b) Effective Rank (DQN)
+(c) Normalized Score (DQN)
+Figure 7: (a) The average fraction of dead neurons, (b) the average effective rank (Kumar et al., 2021), and (c) the baseline-normalized score when training DQN in the Atari domain for 10M iterations (40M frames).
+Similar to the well-known dying ReLU problem, hyperbolic tangent and sigmoid activations also exhibit strong dying neuron behavior.
 A Hadamard product of hyperbolic tangents reduces dead neurons in $z_{t}$ and subsequently increases the effective rank of the representation.
 Mitigating dying Hyperbolic Tangents
-As mentioned in Section 3 and defined in Eq.
+As defined in Eq.
 1, the number of dead neurons is equal to the amount of neurons that display the same saturated output for any given observation $s_{t}$.
-The amount of dead neurons over time when training on Atari can be seen in Fig.
-6(a).
+Over 8 Atari environments, the average amount of dead neurons during training can be seen in Fig.7(a).
 For the ReLU activation, around 60% of the neurons in the representation $z_{t}$ die during training, while for the sigmoid and hyperbolic tangent activation this number is around 40%.
@@
 We credit this to the inherent ability of a Hadamard product of hyperbolic tangents to minimize long-term activation saturation, as explained in Section 4.
-Quantitative results of dead neurons can be found in Table 2, which is correlated with earlier predictions from Table 1.
-Table 2: Dying neuron fractions with and without a Hadamard representation (HR).
-Activation Function Dead Neuron Fraction Dead Neuron Fraction (HR) Difference
-Hyperbolic Tangent 0.39 0.30 -23%
+Quantitative results of dead neurons can be found in Table 2, which tends to confirm the predictions from Table 1.
+As compared to the Hadamard representation, layer normalization notably reduces dead dead hyperbolic tangents to an even lower value.
+Interestingly, the effects of layer normalization show that even if dead neurons are reduced more strongly, it does not always coincide with a higher effective rank (see Fig.
+7(b).
+On the other hand, the Hadamard representation does seem to have a positive effect on both metrics.
+Table 2: Average dying neuron fractions in 8 Atari Games with and without a Hadamard representation (HR).
+Activation Without HR With HR $\Delta$
+Tanh 0.39 0.30 -23%
 Sigmoid 0.44 0.45 +2%
@@
 We additionally investigate the effective rank (Kumar et al., 2021) of the representation $z_{t}$ during training, which can be seen in Fig.
-6(b).
+7(b).
 As observed by Gulcehre et al.
@@
 Furthermore, similar to the results in our supervised learning experiments (see Fig.
-2), using a Hadamard representation with hyperbolic tangents significantly improves the effective rank, which is strongly correlated to a network’s ‘expressivity’.
-Employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected.
-Analyzing Performance in Atari
-The influence of a Hadamard representation on downstream performance can be seen in Fig.
-7.
+2), using a Hadamard representation with hyperbolic tangents significantly improves said effective rank, which is strongly correlated to a network’s ‘expressivity’.
+Alternatively, we found that employing a Hadamard representation with ReLU activations significantly decreased the effective rank of the representation, as was expected from earlier predictions in Table 1.
+More details on the ReLU Hadamard representation, dying neuron calculations and effective rank calculations can be found in Appendices C.4, B.1 and B.2, respectively.
+Performance in Atari
+The influence of a Hadamard representation on downstream performance is visualized in Fig.
+7(c).
 Correlating with the reduction in dying neurons and an increase in effective rank, a significant improvement over the standard hyperbolic tangent baseline is obtained, as well as an improvement over the default ReLU baseline.
@@
 Fig.
-7(a) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
-(a) Normalized Score
-(b) Ablations
-Figure 7: Normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames).
-In (a), the Tanh (HR) significantly outperforms the baseline Tanh.
-The recently proposed Rational Activation (Delfosse et al., 2024) seems to have comparable performance to the ReLU in the 40M frame setting.
-In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced $\alpha=5e^{-5}$ on hyperbolic tangent activations are shown.
+7(c) shows that, although the Rational activation seems to be a stable learnable activation, it remains comparable to the ReLU.
+(a) Ablations (DQN)
+(b) Ablations (DQN)
+(c) Performance (PPO)
+(d) Effective Rank (PPO)
+Figure 8: Baseline-normalized performance with the standard deviation over the means in the Atari domain, after training DQN for 10M iterations (40M Frames).
+In (a), Tanh (HR) significantly outperforms Tanh.
+The Rational Activation (Delfosse et al., 2024) is comparable to ReLU within 40M frames.
+In (b), several ablations including layer normalization (Ba et al., 2016) and an increased dimension of $z_{t}\in\mathbb{R}^{1024}$ with a reduced learning rate $\alpha=5\mathrm{e}{-5}$ are shown.
+In b), Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
+Normalized performance (c) and effective rank (d) when training the PPO algorithm (Schulman et al., 2017).
 Note that no hyperparameter changes are done for the Hadamard representation.
-Figure 8: Additional ablations run on DQN in the Atari domain. Tanh (+) represents an addition rather than a Hadamard product, and Tanh (2HR) uses a triple Hadamard product.
 Examining further ablations in Fig.
-8 shows that using a sophisticated piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance.
+8(a-b) shows that using another piece-wise linear function such as the SELU activation (Klambauer et al., 2017) or using an addition rather than a product of hyperbolic tangents seems detrimental to performance.
 Furthermore, taking a product of 3 hyperbolic tangents (2HR) also appears to enhance performance, though there seems to be a negative effect in the early stages of training as compared to using a single Hadamard product.
@@
 Additional experiments combining the ReLU activation with a Hadamard representation can be found in Appendix C.4.
+The Immunity of the ReLU
+Interestingly, it seems that the ReLU activation’s performance is much less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations.
+As also indicated by Teney et al.
+(2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood.
+However, we have offered a partial explanation for this resilience of the ReLU in Section 4: ReLU saturation represents network pruning while hyperbolic tangent saturation leads to biases in subsequent layers.
 Evaluating Hadamard representations on PPO
 To evaluate results on a policy-based algorithm, additional experiments are run using the PPO algorithm (Schulman et al., 2017).
-For PPO, the internal architectural difference with DQN is that the compressed representation $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients (see Appendix A.2).
-The results can be found in Fig.
-9.
-Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and the best performance.
-The Tanh (HR) and the ReLU also seem to have consistent strong performance across both algorithms, whereas the pure hyperbolic tangent and sigmoid activations can be relatively unreliable across algorithms.
-(a) Normalized Score
-(b) Effective Rank
-Figure 9: Normalized performance (a) and effective rank (b) in the Atari domain for 10M iterations (40M Frames) when training the PPO algorithm (Schulman et al., 2017).
+For PPO, the internal architectural difference with DQN and PQN is that the final hidden layer $z_{t}$ precedes both a critic and an actor network, and thus receives policy and value gradients.
+After training for 40M frames, the performance and effective rank over time can be found in Fig.
+8(c-d).
+Similarly to the results on DQN, the Tanh (HR) exhibits the highest effective rank and good performance.
+The Tanh (HR) and the ReLU seem to have consistent strong performance, whereas the pure hyperbolic tangent and sigmoid activations can be unreliable across algorithms.
+This is coherent with the fact that continuously differentiable activations are generally not favored over the ReLU (Teney et al., 2024).
 Evaluating Hadamard representations on PQN
-To more reliably test the performance gains one can achieve using Hadamard Representations, we evaluate their effects on the Parallelized Q-Network (PQN) in the full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
+To test the performance of the Hadamard representation over a broader selection of environments, we evaluate the recent Parallelized Q-Network (PQN) algorithm in the nearly full Atari suite without hard-exploration environments, for a total of 51 games with 5 seeds per game.
 Figure 9: Median Human-Normalized scores on PQN in 51 Atari games for 5 seeds over 40M frames.
-Labels represent MLP / CNN activations.
-Employing a Hadamard Representation in the Convolutional layers as well as the MLP layer provides a significant performance improvement over the recent PQN baseline, as well as a CReLU (Abbas et al., 2023) implementation of PQN.
-PQN has recently emerged as a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024).
+Labels represent encoder activations.
+Employing a Hadamard representation provides a significant performance improvement over the recent PQN baseline (displayed as ReLU), the CReLU (Abbas et al., 2023) and ReDo ($\tau=0.025$) (Sokar et al., 2023).
+Notably, the improvement over a conventionally parameterized hyperbolic tangent activation is more than 100%.
+PQN is a vectorized-friendly version of DQN, reporting both a speed up from DQN as well as better convergence (Gallici et al., 2024).
 We run tests implementing the Hadamard representation in all hidden layers.
 The median human-normalized scores after training for 40M frames are shown in Fig.
-10.
-We can clearly see a similar trend for the 51-game Atari suite, where using a Hadamard representation for all layers significantly increases the median performance as compared to the baseline PQN and CReLU (Abbas et al., 2023)
-Hyperbolic Tangents turn Weights into Biases
-Interestingly, it seems that the ReLU activations’ performance is less correlated to its low effective rank and high amount of dying neurons than the continuously differentiable activations.
-As also indicated by Teney et al.
-(2024), the exact reasons for the broad success of the ReLU activations are not yet fully understood.
-However, in the case of dying neurons, we aim to shed some light on this phenomenon.
-When dying neurons occur in ReLU-activated layers, it basically prunes these neurons and the associated weights to the next layer.
-However, in hyperbolic tangent activated layers, dying neurons lead to an unintended phenomenon where weights associated with dead neurons effectively become biases.
-Theorem 4.1.
-When any set of neurons $\alpha^{j}$ in a hidden layer $z^{j}$ collapses into nonzero values, the output to the next layer effectively changes from ($A^{j}z^{j}+B^{j}$) to ($A_{-*}^{j}z_{i-*}^{j}+B^{j+1}$ + $A_{*}^{j}z_{i*}^{j}$), where $A_{-*}^{j}z_{-*}^{j}$ represent the active neurons multiplied by their corresponding forward-connected weights and $A_{*}^{j}z_{i*}^{j}~{}=~{}B_{*}^{j+1}$ represent the dead neurons, inducing a hidden bias.
-Proof.
-Take a set of neurons $\alpha^{j}_{i}$ and forward connected weights $\bm{w}^{j}_{\alpha_{i}}$ in layer $z^{j}$. The influence of these neurons on the next layer $z^{j+1}$ is calculated as:
-$z^{j+1}=\sum_{i}\alpha^{j}_{i}\bm{w}^{j}_{\alpha_{i}}+B^{j+1}.$ (2)
-If the set of neurons dies and collapses into 0 ($\alpha^{j}_{i}=0~{}\forall~{}s\in\mathcal{S}$), which occurs when using ReLU activations, the influence on the next layer becomes 0, representing basic pruning:
-$\sum_{i}0\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=0\quad\forall s\in\mathcal{S}.$ (3)
-Where $\bm{\bar{w}}^{j}_{\alpha_{i}}$ is the set of weights connected to the dying neurons.
-However, for a hyperbolic tangent activation, if the set of neurons saturate into either -1 or 1 ($\alpha^{j}_{i}=\{1,-1\}\forall s\in\mathcal{S}$), the output is
-$\sum_{i}\{1,-1\}\cdot\bm{\bar{w}}^{j}_{\alpha_{i}}=B_{*}^{j+1}\quad\forall s\in\mathcal{S}.$ (4)
-As a result, the weight vector $\bm{\bar{w}}^{j}_{\alpha_{i}}$ corresponding to the dying activations only influences a bias $B^{j+1}_{*}$ on the resulting hidden layer.
-Note that the bias $B^{j+1}_{*}$ is constant for any input observation.
-∎
-This emergent bias can hinder the optimization process by introducing unintended fixed contributions to a networks’ next hidden layer or Q-values, reducing the flexibility of the network’s representations and potentially reducing its performance in RL or supervised learning.
-5 Related Work
-Network Capacity in RL.
-Liu et al.
-(2019) investigated the need for sparse representations in the continuous control domain.
-Gulcehre et al.
-(2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues.
-Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022).
-Other work by Lyle et al.
-(2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays.
-Further research by Nikishin et al.
-(2022) used network resets to counteract the primacy bias and Sokar et al.
-(2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting.
-Nikishin et al.
-(2023) further studied plasticity injection for long-term training and Delfosse et al.
-(2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity.
-Concurrent work by Dohare et al.
-(2024) used continual back-propagation to further alleviate plasticity loss.
-In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024).
-This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
-Network Architecture in RL
-The origin of network optimization problems with hyperbolic tangents and sigmoids were empirically investigated by Glorot & Bengio (2010), where, according to the authors, a lot of mystery still surrounds the subject.
-Work by Srivastava et al.
-(2015) in supervised learning first looked at the idea of using products of hidden layers together with a ’gate’ that determined the amount of information flow (Hochreiter & Schmidhuber, 1997).
-Using these ideas, the Resnet was invented (He et al., 2016) and also showed strong performance in combination with RL (Espeholt et al., 2018).
-Further work by Henderson et al.
-(2018) showed differences in RL performances over different network architectures and nonlinear activations.
-Work by Abbas et al.
-(2023) successfully applied ReLU concatenation (Shang et al., 2016) to improve continual learning while keeping a similar performance when training from scratch.
-Finally, recent work by Grooten et al.
-(2024) investigated raw pixel masking for distractions in RL using a parallel CNN input layer.
+9.
+We can clearly see a similar trend for the 51-game Atari suite.
+Using a Hadamard representation significantly increases the median performance as compared to the baseline PQN, the CReLU activation (Abbas et al., 2023), and the ReDo + ReLU algorithm (Sokar et al., 2023).
+Additionally, preliminary modified applications of ReDo to Tanh did not prove to be effective in PQN.
+However, we believe that future work in this direction would be worth pursuing.
+6 Limitations
+Using a Hadamard representation increases neural network parameters as it doubles the incoming weights connected to a hidden layer.
+However, in Fig.
+9, a comparison against the CReLU is shown, which also nearly doubles the network’s parameters.
+Furthermore, recent work shows that simply scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024).
+Another limitation is that, due to the vast amount of activation functions and the theoretical sound candidacy of the hyperbolic tangent, our research focused on the combination of hyperbolic tangents.
+Further research into Hadamard representations could find more novel activation combinations.
+A hyperparameter search will also give more insights into the strengths of the HR, as all baselines are specifically tuned for the ReLU.
+Finally, integration of a HR into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al.
+(2021) would be interesting.
+Table 3: Parameter count across different activations
+Activation Total Parameters
+Tanh (Width2absent2$\cdot$ 2⋅ 2) 6,717,376
+Tanh (HR) 3,373,382
+CReLU 3,292,838
 7 Conclusions and Discussion
-This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem.
-Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer.
-We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness.
-In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
-Finally, we have given insights into the ability of the ReLU activation to be less affected the symptoms of dying neurons, as opposed to non-zero saturating activations such as the hyperbolic tangent.
+This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem, but with more acute downstream effects.
+To alleviate this problem and make better use of these functions’ properties, a novel representation architecture called the Hadamard representation (HR) is proposed.
+The HR augments the hidden layers in a network by taking the Hadamard product with parallel, independently parameterized activation layers.
+We further analyzed and empirically showed that applying a Hadamard representation to hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness.
+In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterization and merely increasing the representation dimension.
 Future work could focus on further identifying the intricacies of the effects that different activation functions have on the resulting representation, in an attempt to push the potential of non piece-wise linear activations in reinforcement learning or even in a supervised learning setting (see Fig.
@@
 (2024) could be promising.
-7 Limitations
-Using a Hadamard representation will double the incoming weights connected to a hidden layer.
-However, in Fig.
-10, a comparison against the CReLU is shown, which equally doubles the network’s parameters.
-Furthermore, recent work shows that a scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024).
-Second, due to the vast amount of activation functions and the theoretical sound candidacy for the hyperbolic tangent, we have mostly focused our research on the combination of hyperbolic tangents.
-Further research into multiplicative representations could find even better activation combinations.
-A hyperparameter search could also give more insights into the strengths of the Hadamard Representation, since the baselines are specifically tuned for the ReLU activation.
-Finally, as this research performs experiments on DQN, PPO and PQN, integration of a Hadamard representation into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al.
-(2021) would be interesting.
-Reproducibility Statement
-Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1.
-The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022).
-The PQN algorithm uses the implementation in the Github repository of Gallici et al.
-(2024).
-Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
 Appendix A Implementation Details
@@
 Number of Test Envs 8 Number of environments used for testing
-Hadamard Implementation
-Constructing a Hadamard representation is a straightforward process that only requires an additional, parallel linear layer.
-Starting from the flattened convolutional features in DQN, the Pytorch pseudocode is defined as follows:
-flattened_features = conv_features.flatten(1)
-representation1 = nn.Tanh(linear1(flattened_features))
-representation2 = nn.Tanh(linear2(flattened_features))
-hadamard = representation1 * representation2
-Q-values = linear_q(hadamard)
-Reinforcement Learning
+A.2 Hadamard Implementation
+Constructing a Hadamard representation is a straightforward process that only requires additional, parallel incoming weights.
+Starting from a hidden layer, the Pytorch pseudocode is defined as follows:
+# hidden = previous hidden layer
+linear1 = nn.Linear(input_dim, output_dim)
+linear2 = nn.Linear(input_dim, output_dim)
+representation1 = nn.Tanh(linear1(hidden))
+representation2 = nn.Tanh(linear2(hidden))
+hadamard_representation = representation1 * representation2
+A.3 Reinforcement Learning
 In DQN, the action $a_{t}$ is chosen following an $\epsilon$-greedy policy.
@@
 This clipped surrogate objective ensures gradual and stable policy updates.
-A.2 PPO Architecture
-Figure 11: A visualisation of the Hadamard representation (HR) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment.
-A parallel linear layer providing an independent representation is integrated, where the Hadamard product of the two parallel representations represents the final representation $z_{t}$.
 Appendix B Kernel Density Estimations
 As discussed in Section 4, we hypothesize that the differences between a hyperbolic tangent with and without an HR are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons.
-We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig.
-12(a).
+We further see this phenomenon when plotting a random selection of neurons from both the mask and the base representation in Fig.
+10(a).
 (a) Tanh (HR) - $5\cdot 10^{6}$ iterations
@@
 The comparison with shallow networks can be found in Fig.
-14(a) and a comparison with deep networks can be found in Fig.
-14(b).
+12(a) and a comparison with deep networks can be found in Fig.
+12(b).
 (a) Comparison of shallow networks for a nonlinear regression task.
@@
 These derivations are empirically validated by the results in Fig.
-16.
+14.
 In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1.
@@
 Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
-C.5 Hyperbolic Tangent Convolutional Activation
-We additionally run tests in DQN on 8 Atari games. The baseline ReLU is used, except for the convolutional activations which are changed to a hyperbolic tangent.
-Figure 18: Baseline normalized performance in the Atari domain for 10M iterations (40M Frames) on DQN.
-It is evident that using non-Hadamard hyperbolic tangents in the convolutional layers can lead to only a fraction of the performance of the ReLU-activated baseline.
 Appendix D Atari
-D.1 DQN Evaluation Details
-We normalize performance with respect to the ReLU baseline in cleanrl  Huang et al.
-(2022).
+D.1 DQN & PPO Evaluation Details
+For DQN and PPO, the Hadamard representation was applied to the final hidden layer of the Nature CNN.
+Furthermore, we have normalized performance with respect to the ReLU baseline on which the experiments were build (Huang et al., 2022).
 The minimum and maximum score of the ReLU baseline are taken for each environment, and the normalized score for each environment is calculated as follows:
@@
 Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig.
-19.
+16.
 Due to taking a subset of the Atari domain in DQN and PPO, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation.
@@
 D.2 Individual Environment Scores
-Figure 20: DQN Performance comparison on the individual Atari Environments.
-Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
-Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations.
-No other games were tested.
-Figure 21: PPO Performance comparison on the individual Atari Environments.
-Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
-Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations.
-No other games were tested.
+Figure 17: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
+Figure 18: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 19: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 20: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 21: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 22: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 23: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 24: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
 Figure 25: PQN Performance comparison on the individual Atari Environments.
-Labels represent MLP / CNN activations.
+Labels represent encoder activations.
 Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.